### v2

# AAAI Press Formatting Instructions for Authors Using — A Guide

Abstract
Clinical decision-making often involves interpreting images (e.g., radiology) for making diagnoses. Retrieving relevant visual information from medical literature and hospital records could enhance diagnostic accuracy. In this paper, we develop a model in which a multimodal retriever is jointly optimized with an LVLM for medical diagnosis, unlike standard RAG where LVLM error signal is not propagated down to the retriever. We show that using only general-purpose backbones, with only lightweight fine-tuning, our model is able to achieve competitive results with medically-pretrained models across clinical multi-label classification and visual question answering tasks. In a novel analysis, we additionally find that in many cases different top retrieved images each lead to different predictions for a given target, and that these cases are empirically challenging for all models, even for non-retrieval models. Our joint retrieval optimization significantly improves these challenging cases over standard RAG. However, oracle analysis reveals that while the correct diagnosis is frequently achievable using one of the top retrieved images, in practice there is a large performance gap from the oracle, and rerankers using frontier LVLMs do not close this gap–leaving ample room for improvement by future methods.Code available at: https://github.com/Nirmaz/JOMED.
Introduction
Figure 1: We jointly optimize a multimodal retriever and a Large Vision-Language Model (LVLM) for medical tasks. Our model achieves competitive results without resource-intensive medical pre-training and significantly improves performance on challenging cases where different retrieved images lead to inconsistent retrieval prediction.
Inferring diagnoses from medical imagery is a fundamental part of clinical decision-making. Large Vision Language Models (LVLMs) have been widely explored for medical diagnosis (Thawakar et al. 2024; Wu et al. 2023; Zhang et al. 2023b; Moor et al. 2023; Li et al. 2023). To improve the performance of LVLMs in the medical field, retrieval augmentation has been adopted and has shown promising results, providing more accurate and also explainable methods (He et al. 2024a; Xia et al. 2024b, a; Wu et al. 2025).
In this work, we explore a lightweight fine-tuning approach for taking a general-purpose LVLM and optimizing it jointly with a general-purpose multimodal retriever. Different from the common retrieval augmented generation (RAG) setup, we perform joint optimization of the LVLM and retriever. This joint optimization teaches the LVLM to utilize retrieved knowledge for making accurate medical predictions while at the same time tuning the multimodal retriever to retrieve information that leads to correct predictions by the LVLM. The retriever model retrieves images as well as captions and clinical reports from both literature and hospital records. Our method involves sequential multimodal training with a dual-head retriever architecture and a retrieval loss that enhances the retriever’s ability to provide more predictive context to the LVLM.
Poorly optimized retrieval mechanisms can mislead models (Yoran et al. 2023; Sun et al. 2024). Figure 2 demonstrates this problem with an ultrasound image that shows no breast cancer. When the retrieval system selects an irrelevant reference image showing cancer features, the model incorrectly identifies cancer in the input image. When we jointly optimize the retriever to find more relevant images and text, the model’s prediction flips to the correct diagnosis.
Using our approach, we are able to achieve competitive results in medical image classification and visual question answering (VQA). Importantly, our focus in this work is not to chase SOTA results but to shed light on several interesting observations. One, that a simple and effective joint LVLM-retriever optimization mechanism can lead to a substantial boost in results over standard multimodal RAG and should thus be more widely adopted for medical diagnosis.
Second, that this mechanism achieves competitive results by using models with no medical pre-training (for neither the LVLM nor the retriever) and only lightweight fine-tuning; this resonates with recent findings showing general-purpose LVLMs can rival their medical counterparts (Jeong et al. 2024) and has potentially important implications considering the resource-intensive nature of pre-training processes.
Third, previous work on joint training in the general domain required large-scale pre-training followed by few-shot task-specific adaptation (Izacard et al. 2023; Hu et al. 2023; Lin et al. 2024). In contrast, our novelty is in conducting joint training directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples). This provides first evidence about the feasibility and utility of lightweight data-efficient joint training directly on downstream tasks as opposed to in the pre-training setting. We are also the first to explore and show the utility of joint training in the multimodal medical domain.
Fourth and importantly, we conduct a novel analysis that finds that our method helps in particular to address a class of errors we term inconsistent retrieval predictions. Inconsistent retrieval predictions are cases in which for a given patient image, each retrieved image leads the model to make different predictions. These cases commonly occur across our experiments and are substantially more difficult for models, both retrieval-augmented and non-retrieval models. This instability with respect to different retrieval candidates leads to high prediction entropy/uncertainty across classes, degrading not only overall RAG results but also their reliability (Lambert et al. 2022).
As we show, our joint training mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases.
We further discover that for a large proportion of these inconsistent cases, at least one retrieved candidate enables the model to predict the correct answer in an oracle setting in which we provide the model with the correct retrieved candidate. This suggests useful information is being retrieved but is often lost among less helpful candidates. To address this, we explored using the powerful multimodal model o3 (OpenAI 2025) as a reranker to identify the most predictive instances. We found it did not close the performance gap to the oracle and was comparable or inferior to our method, leaving significant room for future improvement.
Our contributions are three-fold:
  • We develop a method where a multimodal retriever and a Large Vision-Language Model (LVLM) are jointly optimized for medical classification and VQA. Unlike standard Retrieval-Augmented Generation (RAG), our method propagates error signal from the LVLM down to the retriever while training on downstream tasks.
  • We show that our method achieves competitive results compared to medically pre-trained models by using only general-purpose backbones. Unlike previous general-domain joint training methods that require pre-training, we conduct joint training directly on downstream tasks with only lightweight data-efficient fine-tuning.
  • In a novel analysis we identify inconsistent retrieval predictions, challenging cases where different retrieved images lead to different predictions for the same query. Our joint optimization method significantly improves performance on these cases when compared to standard RAG.
Figure 2: Joint training impacts inconsistent retrieval predictions. After joint training, retrieved images show greater alignment with query image labels. In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after joint training, the retrieved image is less directly related to the wrong label (c). For Retina, joint training changed retrieval from a different modality (e) to be more similar to the query (d). In VQA-RAD, retrieval shifted from an unrelated medical condition (h) to an image depicting the condition of the query image.
Related Work
Retrieval Augmented Joint Training has been explored primarily in unimodal (text only) NLP with work such as ATLAS (Izacard et al. 2023), which conducted large-scale pretraining of both reader and retriever models and evaluated in zero-shot and few-shot scenarios in the general domain. REVEAL (Hu et al. 2023) extended this to the multimodal setting, achieving state-of-the-art results in visual tasks in the general domain with extensive pre-training. (Shi et al. 2023) proposed training textual retrievers based on reader performance, influencing subsequent joint-training methods (Lin et al. 2024). introduced another joint training variant and proposed sequential training: first training the reader with retrieval augmentation, then training the retriever. Siriwardhana et al. (2023) demonstrated joint training’s effectiveness for domain adaptation. Our work differs by being the first to explore joint training optimized directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training. Our extensive novel analyses are the first to shed light on inconsistent predictions and how joint training on downstream tasks helps address them.
Multimodal Retrieval Augmentation in Medical Applications. Multimodal retrieval augmentation began with encoder-based architectures (Yuan et al. 2023), which integrated retrieved text and images with query modalities. With Large Vision-Language Models (LVLMs), (He et al. 2024a) proposed retrieving relevant labeled examples during inference. (Xia et al. 2024b) addressed LVLMs’ misalignment with medical knowledge using retrieval-augmented generation to fetch similar medical reports, employing contrastive learning and context selection strategies. (Xia et al. 2024a) proposed domain-aware retrieval for diverse medical data sources, while (Wu et al. 2025) combined retrieval augmentation with knowledge graphs to reduce hallucination.
Method
Figure 3: Jomed combines a general-purpose VLLM with a general-purpose multi-modal retriever in a two-phase approach. First, with a frozen retriever, we augment training by prepending retrieved (image, caption/report) pairs to original queries, training the LVLM on these augmented prompts. Second, with the LVLM frozen, we sequentially train the dual-head multi-modal retriever using Sharp Perplexity Distillation loss (KL-divergence between the distribution derived from the model logits and the distribution derived from the retrieved candidates’ similarity).
Overview. In this section, we present our methodology, named Jomed (Joint Optimization of Multimodal Retrieval for Enhanced Medical Diagnosis). Our task involves classifying medical images and medical visual question answering. Jomed consists of two main components: a multimodal retriever and a reader. We are given an input patient image and a diagnostic question. The multimodal retriever is tasked with identifying relevant medical knowledge from either literature or hospital records—in the form of images and associated captions or hospital reports that provide predictive information. The reader then analyzes the retrieved candidates along with the diagnostic question and generates an answer. We jointly optimize the reader to better utilize retrieved candidates, and the multimodal retriever to improve its ability to retrieve informative candidates for the reader. Figure 3 illustrates our method. The multimodal retriever has two retrieval heads, textual and visual, for finding candidates in an index of image-caption pairs derived from literature and hospital records. Given an image query, these retrievers find candidates that are then passed to the reader.
Unlike previous work (He et al. 2024a; Xia et al. 2024b, a) our method does not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip (Xiao, Mastrapas, and Wang 2024). The selected LVLMs can process multiple images and texts, allowing us to incorporate both the retrieved image and the retrieved caption/report into the model context. We also compare our results to Med-Flamingo (Moor et al. 2023). MedFlamingo is a medical LVLM which is also able to process multiple images, unlike other medical LVLM available at the time of conducting our experiments. Finally, we also explore the utility of o3 (OpenAI 2025) as a reranker. We now elaborate on the details.
Reader Retrieval Augmentation Fine-Tuning
We first fine-tune the reader with retrieval augmentation. This step aims to achieve two main goals: improving the reader’s performance on the dataset and teaching the reader to effectively utilize retrieved (image, text) pairs in context.
For an image, question tuple $(i_{d},q_{d})$ $d\in\{1,\ldots,D\}$ from dataset $D$ , we retrieve $K$ relevant tuples of images and texts (reports or captions) $(i_{k},t_{k})$ $k\in\{1,\ldots,K\}$ using our multimodal retriever module. These retrieved pairs are prepended to the image-question pair $(i_{d},q_{d})$ to create augmented inputs. The augmented input consists of the retrieved context followed by the target question-image pair: $(i_{k},t_{k}),(i_{d},q_{d})$ $k\in\{1,\ldots,K\}$. To the best of our knowledge, we are the first to use both the retrieved image and caption/report in context within medical RAG for LVLMs. We use supervised fine-tuning on the augmented input, minimizing the negative log-likelihood to predict the answer $a_{d}$:
$\mathcal{L}(\theta)=-\sum_{d=1}^{D}\sum_{k=1}^{K}\log p_{\theta}(a_{d}|(i_{k},t_{k})\circ(i_{d},q_{d}))$
where $\theta$ represents the parameters of the LVLM, and $\circ$ denotes the concatenation operation in the context window.
Sequential Multimodal Retriever Fine-Tuning
Our dual-head multimodal retriever has a text-based head for retrieving relevant (image, text) pairs based on textual similarity to the query, and an image-based head for retrieving pairs based on visual similarity. To train this multimodal retriever, we measure how each retrieved (image, text) pair contributes to the LVLM’s performance, optimizing the retriever to prioritize high-contribution candidates.
We adopt a sequential training strategy, first optimizing the text retrieval head followed by the image retrieval head. During this process, the reader model remains frozen, focusing solely on improving the retriever’s embedding space. We formulate a Sharp Perplexity Loss as a simple variant of the perplexity distillation loss (Izacard et al. 2023). Similarly to Izacard et al. (2023) we minimize the KL-divergence between the LVLM’s posterior distribution over retrieved pairs and the retriever’s distribution, however we do so directly on downstream tasks with only lightweight fine-tuning in the data-efficient regime without pre-training.
The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved (image, text) pair $(\mathbf{i}_{k},\mathbf{t}_{k})$ and a query image-question pair $(\mathbf{i},\mathbf{q})$:
$p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q})),$
where $p_{\text{LVLM}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q}))$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
We sharpen the LVLM’s output distribution by restricting it to only the relevant class tokens from the benchmark’s possible answers. Specifically, we extract the model’s logits, select only the relevant tokens for each class, and apply a softmax to obtain a more discriminative distribution (for VQA tasks, we use yes/no tokens and skip open questions.). We denote this class-restricted distribution as $p_{\text{LVLM}_{C}}$ where $C=\{c_{1},c_{2},c_{3},\ldots\}$ is the set of chosen class tokens. The normalized posterior $p_{k}$ is formulated as:
$p_{k}=\frac{\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q})))}{\sum_{i=1}^{K}\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid(\mathbf{i}_{i},\mathbf{t}_{i})\circ(\mathbf{i},\mathbf{q})))},$
where $K$ is the total number of retrieved pairs.
The retriever’s distribution is defined by:
$p_{\text{RETR}}((\mathbf{i},\mathbf{t})\mid(\mathbf{i},\mathbf{q}))=\frac{\exp(s((\mathbf{i},\mathbf{t}),(\mathbf{i},\mathbf{q}))/\tau)}{\sum_{k=1}^{K}\exp(s((\mathbf{i}_{k},\mathbf{t}_{k}),(\mathbf{i},\mathbf{q}))/\tau)},$
where $s((\mathbf{i},\mathbf{t}),(\mathbf{i},\mathbf{q}))$ is the similarity score between the pair and the query, and temperature $\tau$ controls distribution sharpness. For the text retrieval head, the score is computed using the dot product between the index caption/report embeddings and the query image embedding, while for the image retrieval head, we use the index image embeddings.
Finally, for our loss we compute the KL-divergence between retriever and reader distributions $\text{KL}(p_{\text{LVLM}_{C}}\|p_{\text{RETR}})$:
$\sum_{k=1}^{K}p_{\text{LVLM}_{C}}((\mathbf{i}_{k},\mathbf{t}_{k}))\log\left(\frac{p_{\text{LVLM}_{C}}((\mathbf{i}_{k},\mathbf{t}_{k}))}{p_{\text{RETR}}((\mathbf{i}_{k},\mathbf{t}_{k}))}\right)$
Inference
Following Shi et al. (2023) for a given (image, question) pair, we retrieve the top-$k$ relevant (image, text) chunks using the image as the query. Each chunk is prepended to the question, and the LVLM computes predictions for the augmented prompts in parallel. The final output probability is:
$p_{\text{LVLM}}(a\mid\mathbf{q})=\sum_{k=1}^{K}p_{\text{LVLM}}(a\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ\mathbf{q})\cdot p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q}),$
where $p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q})$ is:
$p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q})=\frac{\exp(s(\mathbf{q},(\mathbf{i}_{k},\mathbf{t}_{k})))}{\sum_{j=1}^{K}\exp(s(\mathbf{q},(\mathbf{i}_{j},\mathbf{t}_{j})))},$
and $s(\mathbf{q},(\mathbf{i}_{k},\mathbf{t}_{k}))$ is the similarity score between the query and the retrieved chunk.
Backbone Model Breast Derma Retina VinDr-PCXR BRSET Mean
ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
Qwen2-vl Reader .83 .77 .68 .27 .60 .44 .48 .08 .34 .23 .59 .36
FT RAG .85 .82 .71 .42 .62 .48 .55 .09 .48 .27 .64 .42
Jomed .87 .84 .76 .50 .65 .50 .57 .14 .49 .37 .67 .47
Jomedoracle .87 .84 .85 .64 .69 .51 .64 .14 .52 .41 .71 .51
Pixtral Reader .82 .77 .75 .52 .55 .44 .48 .08 .44 .33 .61 .43
FT RAG .88 .85 .79 .60 .57 .47 .49 .09 .47 .33 .64 .47
Jomed .90 .87 .80 .62 .60 .51 .56 .14 .51 .37 .67 .50
Jomedoracle .93 .90 .83 .67 .63 .54 .67 .15 .57 .41 .73 .53
Backbone Model VQA-RAD SLAKE PathVQA Mean
Closed Open Closed Open Closed Open Closed Open
Qwen2-vl Reader .73 .41 .84 .80 .87 .25 .81 .49
FT RAG .76 .45 .88 .81 .91 .33 .85 .53
Jomed .78 .47 .90 .83 .93 .37 .87 .56
Jomedoracle .86 - .92 - .95 - .91 -
Pixtral Reader .72 .38 .83 .80 .87 .26 .81 .48
FT RAG .74 .41 .88 .81 .88 .31 .83 .51
Jomed 76 .43 .90 .82 .90 .35 .85 .53
Jomedoracle .88 - .92 - .95 - .92 -
Table 1: Results for classification (a) and visual question answering (b). Joint training consistently leads to the best results, with up to a 10-point increase in F1 score compared to fine-tuned (FT) RAG for multi-label classification. For visual question answering, we evaluate closed questions using the exact match metric and open questions using token recall. Using an oracle reranker demonstrates that even larger improvements are achievable with Jomed’s top-retrieved images in some datasets. Note that in the open question scenario, we do not evaluate the oracle because there is no clear discriminative boundary between different answers to determine the correct one, as there is in classification and closed question answering tasks. Using an oracle reranker shows that larger improvement is achievable with Jomed’s top retrieved images in some datasets.
Experiments
Medical
Pretrained
Size
Model Breast Derma
177K MedVInT-TE .88 .78
177K MedVInT-TD .90 .80
177K MedDrInternVL .72 -
177K MedDr + RADInternVL .88 -
255K GSCoInternVL .93 -
0 Jomed Qwen2-vl .87 .76
0 JomedPixtral .90 .80
Table 2: Accuracy of Jomed and prior reported results of medical pre-trained models for multiclass classification.
Medical
Pretrained
Size
Model VinDr-PCXR BRSET
255K MedDrInternVL .08 .08
255K GSCoInternVL .09 .33
0 Jomed Qwen2-vl .14 .37
0 JomedPixtral .14 .37
Table 3: F1 Comparison of Jomed to prior reported results of medical pre-trained models for multi-label classification.
Medical
Pretrained
Size
Model SLAKE PathVQA
Closed Open Closed Open Open
600K BiomedGPT-B .90 .85 .88 .28
600K LLaVA-MedLLaVA .83 .85 .91 .38
600K LLaVA-MedVicuna .85 .83 .92 .39
600K LLaVA-MedBioMedCLIP .87 .87 .91 .39
0 JomedQwen2-vl .90 .83 (.84) .93 .37 (.37)
0 JomedPixtral .90 .82 (.82) .90 .35 (.35)
Table 4: Comparison of Jomed to medical pre-trained models for visual question answering. We report the token recall metric, reported for LLaVA-Med variants and token F1 reported for BiomedGPT (in red).
Experimental Setup
Our datasets include real-world hospital datasets (BRSET (Nakayama et al. 2024) and VinDr‑PCXR (Pham, Tran, and Nguyen 2022)) alongside a variety of classification and visual question answering datasets. We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA). Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive. Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed (Casey 2022). Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study. Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge to compensate for sparse training samples.
We consider three types of classification tasks: binary, multi-class, and multi-label. These include BreastMNIST (nickname: Breast; breast ultrasound imaging, binary) (Al-Dhabyani et al. 2020), DermaMNIST (nickname: Derma; pigmented skin lesion images, multi-class) (Tschandl, Rosendahl, and Kittler 2018), RetinaMNIST (nickname: Retina; retinal fundus images with multi-label disease annotations) (Liu et al. 2022), and two challenging real-world multi-label datasets: VinDr-PCXR (a chest X-ray dataset with 15 labels) (Pham, Tran, and Nguyen 2022) and BRSET (an ophthalmology dataset with 14 labels) (Nakayama et al. 2024). For visual question answering, we use widely adopted benchmarks: VQA‑RAD (Lau et al. 2018), SLAKE‑English (Liu et al. 2021), and PathVQA (He et al. 2020). Full details are provided in Appendix A.
Retrieval augmentation is supported by an external index constructed from PubMed and medical records: PMC‑OA (Lin et al. 2023), MIMIC‑CXR (Johnson et al. 2019), and ROCO (Rückert et al. 2024). Full descriptions are available in Appendix A. Our experiments leverage two backbone models—Pixtral (12B) (Agrawal et al. 2024) and Qwen2‑vl (7B parameters) (Wang et al. 2024)—with jina‑clip‑v1 (Xiao, Mastrapas, and Wang 2024) serving as the general-purpose retriever’s visual head, embedding both the texts and images of the retrieval index. All models were trained in a low-resource setting, with all runs performed on a single L40s GPU. More details in Appendix B.
For comparison to medically pre‑trained LVLMs, we include competitive baseline models: BiomedGPT (Luo et al. 2023); three LLaVA-Med (Li et al. 2023) variants; two LVLM variants based on PMC CLIP (MedVInT-TE and MedVInT-TD) (Zhang et al. 2023b); and three variants based on InternVL (Chen et al. 2024): MedDr, MedDr + RAD (He et al. 2024a), and GSCo (He et al. 2024b). Full details are in Appendix C.
Results
Table 1(a) and Table 1(b) demonstrate the effectiveness of Jomed for two backbone models, Pixtral and Qwen2-vl, across five medical imaging classification benchmarks and three visual question answering benchmarks. Jomed achieved superior results compared to both the standard fine-tuned multimodal RAG baseline (trained only with retrieval augmentation fine-tuning, as detailed in Section Reader Retrieval Augmentation Fine-Tuning) and the Reader baseline (LVLM trained without retrieval), for both backbones and all benchmarks.
In Table 1(a), for classification tasks, Jomed with the Qwen2-vl backbone achieved a mean improvement of 0.05 in F1 and 0.03 in ACC over fine-tuned multimodal RAG, with F1 gains of 0.10 and 0.08 in two datasets. Jomed with Pixtral achieved a mean improvement of 0.03 in F1 and ACC. In Table 1(b), for VQA tasks, Jomed with Pixtral and Qwen2-vl achieved a mean improvement of 0.02 in ACC (Exact Match) for closed question answering. For open question answering, Jomed with Qwen2-vl achieved a mean improvement of 0.03 in token recall, while Jomed with Pixtral achieved an improvement of 0.02.
In addition, we evaluated our model in an oracle scenario. Instead of the final fusion of logits (Section Inference), we applied an oracle that, given the model responses for different retrieved images, selects the correct answer if it exists. Our findings in Tables 1(a) and 1(b) show that the correct answer exists in a large proportion of cases, yielding superior results. This shows that in these cases, Jomed is able to surface in its top retrieved images an image which could lead to a correct prediction, however the information gets lost in the process of prediction fusion with other retrieved images. Simply taking the label with highest confidence or label with highest average confidence across retrieved images, underperforms. This motivates us to explore the use of the o3 multimodal reasoning model to detect this image (see below).
Benchmarking versus pretrained LVLMs for classification. In Tables 2 and 3, we include previously reported results for several leading medically pretrained LVLM methods (detailed in Section Experimental Setup). Jomed with both backbones delivered competitive performance relative to models that underwent extensive pretraining. For multi-label classification, Jomed outperformed the pre-trained models. For binary and multi-class classification tasks, Jomed with Pixtral matched or exceeded all models except GSCo, which achieved 0.03 higher ACC on the Breast dataset. Overall, our models performed strongly across four medical benchmarks compared to medical pretrained models. We are not aware of any medical LVLM evaluated on Retina. To our knowledge, no other LVLMs have been evaluated on our benchmarks. Additionally, in Table 2 we report only ACC, as F1 scores were not reported for the compared models. Similarly, in Table 3, we report only F1 scores. We also evaluated a medical baseline, Med-Flamingo, with a pretrained medical retriever (BiomedCLIP (Johnson, Douze, and Jégou 2019)). We observed that Med-Flamingo generally underperformed, and showed limited improvement from utilizing the retrieved context to enhance performance, which we suspect may be attributed to its relatively older LLaMA backbone (Touvron et al. 2023). We further try BiomedCLIP as a retriever with our general-purpose LVLMs, which showed improvements over FT RAG, with a similar pattern to the general-purpose retriever. Full results in Appendix D.
Benchmarking versus pretrained LVLMs for visual question answering. In Table 4, we also compare our approach against medical pretrained LVLM methods (methods detailed in Section Experimental Setup) for VQA. Here too, Jomed delivered competitive performance relative to extensively pretrained models. For closed question tasks, Jomed with Pixtral and Qwen2-vl was superior to or matched current models. For open-ended question answering, Jomed with Qwen2-vl backbone matched LLaVA-MedVicuna on the SLAKE dataset and achieved slightly lower results for PathVQA, showing a 0.01 - 0.02 token recall gap compared to pretrained LVLMs. Note that we do not compare our method on VQA-RAD, since our approach is trained on an internal split of training, validation, and test sets, whereas VQA-RAD provides only official training and test splits.
Backbone Model Breast Derma Retina VinDr BRSET VQA-RAD SLAKE Mean Mean
ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
Closed
Acc
Closed
Acc
ACC F1
Qwen2-vl Inconsistent Reader .40 .40 .50 .25 .35 .31 .41 .07 .33 .21 .43 .70 .40 .25
RAG .40 .40 .52 .36 .40 .28 .46 .07 .49 .25 .33 .84 .45 .27
Jomed .80 .80 .62 .46 .52 .29 .48 .16 .50 .36 .47 .86 .58 .41
Consistent Reader .84 .77 .83 .29 .63 .45 .56 .09 .34 .22 .77 .85 .64 .36
RAG .86 .82 .90 .49 .66 .50 .66 .10 .43 .23 .82 .89 .70 .43
Jomed .88 .84 .91 .50 .66 .55 .66 .10 .41 .37 .82 .91 .70 .47
Pixtral Inconsistent Reader .60 .59 .45 .33 .40 .32 .38 .07 .46 .28 .66 .64 .46 .32
RAG .72 .71 .44 .45 .35 .37 .38 .08 .39 .26 .72 .65 .46 .37
Jomed .84 .84 .50 .50 .42 .46 .49 .13 .45 .35 .77 .68 .54 .46
Consistent Reader .86 .80 .80 .56 .56 .44 .66 .08 .61 .37 .74 .85 .70 .45
RAG .91 .87 .84 .67 .62 .50 .71 .08 .64 .38 .75 .91 .74 .50
Jomed .91 .87 .85 .67 .62 .53 .71 .10 .64 .42 .75 .91 .75 .52
Table 5: Analysis of joint training effect: for inconsistent retrieval predictions, we can see that applying joint training boosts performance significantly for all datasets. For consistent retrieval predictions, the results are slightly improved. Note that inconsistent retrievals are challenging to all models, even non-retrieval ones.
Model Breast Derma Retina VinDr BRSET VQARAD SLAKE
Qwen2-vl 3% 51% 15% 50% 93% 12% 16%
Pixtral 16% 13% 16% 66% 70% 7% 8%
Table 6: Proportion of cases with inconsistent predictions, per dataset, for Jomed based on Qwen2-vl and Pixtral.
Analysis
Jomed boosts empirically challenging cases. We observed that the performance gap is substantially larger for predictions classified as inconsistent retrieval predictions (the proportion of inconsistent retrieval predictions is reported in Table 6). We define an inconsistent retrieval prediction as one in which the model predicts different labels for retrieved candidates given the same target query (see Figure 1). Specifically, if the model’s predictions vary across retrieved candidates, the instance is categorized as an inconsistent retrieval prediction. Conversely, if the model consistently predicts the same label for all retrieved candidates, it is classified as a consistent retrieval prediction.
Both the consistent and inconsistent retrieval prediction sets are established after the initial reader training stage, and their performance is assessed following joint multimodal retriever fine-tuning. Detailed results are presented in Table 5. Overall, Jomed improves performance on inconsistent retrieval predictions by +0.12 in accuracy and +0.13 in F1 score when using the Qwen2‑vl backbone, and by +0.09 in both accuracy and F1 score with the Pixtral backbone, while also offering a slight improvement on the consistent.
We also observed that inconsistent retrieval predictions are empirically more challenging for all models (both with and without retrieval). There is approximately a 10–20 point performance gap between the inconsistent and consistent.
Performance with o3 as reranker. We have observed that results for inconsistent predictions remain relatively low (Table 5), even after applying joint training (Table 5). However, in the oracle analysis reported above, we found that Jomed can sometimes retrieve images that lead to correct predictions, together with other images that lead to wrong predictions. We thus explore whether a state-of-the-art multimodal model, when used as an optional reranker, could close the gap toward oracle performance for inconsistent retrieval predictions. Specifically, we investigated whether the o3 reasoning model (OpenAI 2025) could effectively select the image from the retrieved set that offers the most predictive information, guiding the model toward correct classification. We provided o3 with an image to classify, along with four retrieved images+captions, and tasked it with identifying the one containing the most predictive information. (detailed prompt in Appendix D)
Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in Jomed. Thus, we were unable to bridge the gap to oracle performance, and conclude that reranking in this setting remains a significant challenge even for a frontier model. We report full results in Appendix D.
Additional analyses and ablations. We include several robustness ablations and analyses in Appendix D. We observed that our model is robust to scenarios where the retrieval is noisy or not given to the models. Further, we observed that the retrieval itself includes predictive knowledge without being given the query. details in Appendix D.
Conclusion
We explored the effectiveness of a multimodal retriever jointly optimized with a general-purpose LVLM. Our lightweight approach yielded substantial improvements, including in challenging cases involving inconsistent retrieval predictions. Our oracle analysis revealed a considerable performance gap between actual results and what was theoretically achievable using the retrieved images, providing a foundation for future research on closing this gap.
Appendix A Dataset
Backbone Retrieval type Breast Derma Retina VQA-RAD Mean
ACC F1 ACC F1 ACC F1 ACC ACC F1
Flip a coin .46 .30 .14 .10 .17 .16 .50 .35 .31
WO Q img .75 .42 .66 .18 .47 .33 .67 .67 .49
Pixtral Reader only .82 .66 .75 .53 .56 .44 .72 .74 .65
WO Retrieval .84 .65 .78 .55 .58 .48 .72 .74 .65
Random Retrieval .84 .69 .78 .56 .56 .45 .70 .74 .65
WO Q img .81 .52 .56 .20 .47 .20 .70 .68 .48
Qwen2-vl Reader only .82 .63 .68 .27 .61 .44 .74 .74 .59
WO Retrieval .86 .70 .70 .33 .62 .42 .76 .76 .62
Random Retrieval .85 .67 .73 .39 .63 .42 .77 .76 .62
Table 7: We evaluate our models in three scenarios: (1) the model is not given the query image and must rely solely on the retrieved context for prediction; (2) the model is not given any retrieval; and (3) the model is given a retrieved context that is randomly chosen and may mislead it. We use two baselines: a random classifier (uniformly selects a class) and a reader-only model (trained without retrieval). We evaluate on Breast, Derma, Retina, and the closed-question subset of VQA-RAD.
Dataset for Evaluation
Our datasets include real-world hospital datasets (BRSET (Nakayama et al. 2024) and VinDr‑PCXR (Pham, Tran, and Nguyen 2022)) alongside a variety of classification and visual question answering datasets. We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA). Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive. Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed (Casey 2022). Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study. Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge to compensate for sparse training samples. For the classification tasks, we used the following datasets:
BreastMNIST (nickname: Breast) (Al-Dhabyani et al. 2020) is a binary classification dataset of breast ultrasound. Following the official split, we use 546 for training, 78 for validation and 156 for testing.
RetinaMNIST (nickname: Retina) (Liu et al. 2022) is a multi-label classification dataset of retina Fundus Camera. Following the official split, we use 1,080 for training, 120 for validation and 400 for testing.
DermaMNIST (nickname: Derma) (Tschandl, Rosendahl, and Kittler 2018) is a multi-class classification dataset of common pigmented skin lesions. Following the official split, we use 7,007 for training, 1,003 for validation and 2,005 for testing. The dataset contains clinical images from 7 different diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis, dermatofibroma, melanoma, melanocytic nevi, and vascular lesions.
VinDr-PCXR (Pham, Tran, and Nguyen 2022) is a large, open pediatric chest X-ray dataset collected in Vietnam (2020–2021) containing 9,125 posteroanterior scans from patients under 10 years old. It provides both lesion-level bounding-box annotations for 36 findings and image-level labels for multi-label of 15 diagnoses, curated by experienced radiologists. We follow the official split, which includes 7,728 training and 1,397 test images, with an additional internal split of the training data into 85% for training and 15% for validation.
BRSET (Nakayama et al. 2024) is a Brazilian multilabel ophthalmological dataset comprising retinal fundus images annotated with 14 distinct pathological findings. The dataset, composed of 16,266 images, presents a challenging real-world scenario for multilabel classification. There is no publicly available split; therefore, we apply an internal split for train, validation, and test sets (70% train, 10% validation, and 20% test).
Breast, Derma, and Retina taken from the large-scale MNIST-like collection of standardized biomedical images, including 12 datasets for 2D and 6 datasets for 3D. We especially used MedMNIST+ (Yang et al. 2023), which is a higher resolution extension of the original MedMNIST. We used the highest resolution of 224 × 224.
For medical question answering, we have three well-known datasets:
VQA-RAD (Lau et al. 2018) is a clinician-annotated visual question answering dataset focused on radiology images (e.g., X-ray, CT, MRI). It pairs each image with both open-ended and yes/no questions. Following the official split, we use 1,753 questions for training and 453 for testing. In addition, we further partition the training set into 85% for training and 15% for validation.
PathVQA (He et al. 2020) is a pathology-focused medical VQA dataset built from textbooks and the PEIR digital library, comprising 4,289 images and 32,632 question–answer pairs spanning both open-ended and yes/no questions. We follow the official split provided by the authors: 19,700 for training, 6,260 for validation, and 6,720 for testing.
SLAKE (Liu et al. 2021) is a bilingual medical VQA dataset. We use the English subset, SLAKE-English, which comprises 642 radiology images and 7.03k English QA pairs. We follow the official train/validation/test split of 4.92k/1.05k/1.06k QA pairs, and include all question types, covering both open-ended and closed-ended formats.
Dataset for Index
For construction of the Index we used three large datasets of (image, text) pairs: PMC-OA (Lin et al. 2023), ROCO (Rückert et al. 2024) and MIMIC-CXR (Johnson et al. 2019):
PMC-OA: is a large-scale dataset that contains 1.65M image-text pairs. The figures and captions from PubMed Central, 2,478,267 available papers are covered.
ROCO: is an image-caption dataset collected from PubMed. It filters out all the compound or non-radiological images, and consists of 81K samples.
MIMIC-CXR: is the largest chest X-ray dataset, containing 377,110 samples (image-report pairs). Each image is paired with a clinical report describing findings from doctors.
Backbone Model Breast Derma Retina
ACC F1 ACC F1 ACC F1
Qwen2-vl Top-1 .50 .60 .32 .25 .38 .32
Top-1 logits .60 .58 .43 .35 .37 .38
o3 .77 .80 .42 .32 .35 .33
o3 multi-image .77 .80 .43 .33 .43 .35
Jomed .77 .80 .57 .43 .51 .47
Pixtral Top-1 .72 .75 .46 .48 .29 .33
Top-1 logits .64 .63 .40 .40 .29 .32
o3 .73 .73 .42 .37 .25 .30
o3 multi-image .73 .73 .43 .42 .34 .37
Jomed .76 .77 .45 .47 .34 .37
Table 8: Study of o3 as a reranker. We evaluate o3’s ability to rerank the inconsistent predictions, selecting which of the retrieved (image, caption) pairs has the most predictive information for the query image. We compared the performance of choosing the highest similarity retrieved (image, caption) pair, choosing the highest confidence (top logits), using only the caption for reranking, and using our current strategy of fusing logits from all predictions. The metrics used are accuracy and F1 score.
Appendix B Implementation Details
Retrieval Augmentation
As described earlier, for each image–question pair, we retrieve $r$ image–text pairs, with $r=4$. We use the input image as the query and Jina-CLIP (Xiao, Mastrapas, and Wang 2024) as the retriever head in our multimodal retriever. The index is constructed with FAISS (Johnson, Douze, and Jégou 2019) over MIMIC-CXR, PMC-OA, and ROCO, which are fully described in Section A in Appendix A. To embed images, we use the Jina-CLIP visual head (the same model used for retrieval); to embed captions/reports, we use the Jina-CLIP text head. The index is stored in float16 due to storage constraints. We also experimented with BiomedCLIP (Zhang et al. 2023a).
Reader Fine-Tuning
We use Pixtral (12B) (Agrawal et al. 2024) and Qwen2-vl (7B) (Wang et al. 2024) as base models. We train the models with and without retrieval augmentation to assess their effect. The retrieval-augmented prompt is described in Appendix E. Note, we also tried Med-Flamingo (9B) (Moor et al. 2023) as a base model.
We fine-tune Pixtral and Qwen2-vl using LLaMA-Factory (Zheng et al. 2024) on a single NVIDIA L40 (48 GB) GPU for 10 epochs. We use a learning rate of $2\times 10^{-5}$ and apply Low-Rank Adaptation (LoRA) (Hu et al. 2021) for parameter-efficient fine-tuning. To all models (including baseline models), we conduct a grid search over batch size (2, 4, 6) and whether to freeze the vision head, selecting the best configuration by validation performance. For Med-Flamingo, we fine-tune using our codebase on a single NVIDIA L40 (48 GB) GPU for 10 epochs. Following the Med-Flamingo paper, the language model and image encoder are frozen, and only the Gated Cross-Attention layers and the Perceiver Resampler are optimized for stable and efficient learning. We use a learning rate of $2\times 10^{-5}$.
Multimodal Retriever Fine-Tuning
We train the retrieval model to fetch relevant context while keeping the reader frozen. We use Jina-CLIP’s visual head as the base model for the multimodal retriever. The model is trained on a single NVIDIA L40 (48 GB) GPU for 100 epochs with a learning rate of $2\times 10^{-5}$, freezing all layers except the last ten. We set the number of retrieved candidates per query to 50. In our experiments, retrieving more candidates further improved retriever performance.
Evaluation Metrics
Classification.
We report Accuracy (ACC) and Macro F1:
$\mathrm{ACC}=\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}[\hat{y}_{n}=y_{n}],\qquad$
$\mathrm{MacroF1}=\frac{1}{C}\sum_{c=1}^{C}\frac{2\,\mathrm{Prec}_{c}\,\mathrm{Rec}_{c}}{\mathrm{Prec}_{c}+\mathrm{Rec}_{c}},$
where $\mathrm{Prec}_{c}$ and $\mathrm{Rec}_{c}$ are precision and recall computed per class $c$, $C$ is the number of classes, and $N$ is the number of examples.
Visual Question Answering (VQA).
For open-ended questions, we compute token-level Recall and F1 between the predicted token set $\hat{A}$ and the reference token set $A$:
$\mathrm{Prec}=\frac{|\hat{A}\cap A|}{|\hat{A}|},\quad\mathrm{Rec}=\frac{|\hat{A}\cap A|}{|A|},\quad\mathrm{F1}=\frac{2\,\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}.$
For closed-ended questions, we use Exact Match.
Backbone
BiomedCLIP
Variant
Breast Derma Retina Vindr BREST VQA-RAD SLAKE PathVQA Mean
ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC ACC ACC ACC F1
Qwen2-vl FT RAG .84 .81 .73 .38 .63 .45 .54 .09 .49 .27 .77 .89 .92 .72 .40
Jomed .87 .82 .75 .40 .65 .50 .58 .15 .50 .39 .79 .91 .92 .74 .45
Pixtral FT RAG .88 .84 .79 .57 .57 .46 .55 .09 .44 .31 .74 .87 .88 .72 .45
Jomed .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
Table 9: Performance comparison of Jomed against FT RAG baseline across medical imaging classification and VQA tasks. Both Qwen2-VL and Pixtral backbones show consistent improvements with Jomed, particularly in F1 scores. Results demonstrate that BiomedCLIP-based retrieval enhances performance similarly to the general-purpose retriever. VQA results are reported for the closed-question subset using exact match accuracy.
MedFlamingo
Variant
Breast Derma Retina VQARAD
ACC F1 ACC F1 ACC F1 ACC
Reader only .82 .77 .67 .54 .57 .40 .65
Jomed .81 .77 .68 .54 .60 .51 .69
Table 10: Med-Flamingo comparison with and without Jomed retrieval augmentation across medical imaging tasks. Jomed achieves best performance on 4 out of 7 metrics, with notable improvements over Reader Only in Retina classification (ACC: 0.57→0.60, F1: 0.40→0.51) and VQA-RAD (0.65→0.69). However, the overall performance gains are more modest than those observed with other backbone architectures, indicating that Med-Flamingo may have limited capacity to leverage additional retrieval context.
Appendix C Baseline Methods
We benchmark against state-of-the-art medical large vision–language models (LVLMs) that underwent large-scale medical pre-training:
BiomedGPT.
An open multimodal generative pre-trained transformer for biomedicine that aligns diverse biological/biomedical modalities with natural language; we include it as a strong domain baseline (Luo et al. 2023).
LLaVA-Med.
A biomedical adaptation of LLaVA trained via curriculum (biomedical figure–caption alignment followed by instruction tuning); we evaluate three commonly used releases as separate baselines (Li et al. 2023).
MedVInT-TE and MedVInT-TD.
Two implementations of Medical Visual Instruction Tuning from PMC-VQA (Zhang et al. 2023b). Both adopt a PMC-CLIP vision encoder (Lin et al. 2023). TE uses an encoder-style text pathway feeding a multimodal decoder, while TD concatenates text tokens with visual features to a decoder-only pathway; we report both as distinct baselines.
InternVL-based baselines: MedDr, MedDr+RAD, and GSCo.
MedDr is a generalist medical VLLM trained on large-scale instruction-style data curated from diagnosis-guided bootstrapping and medical image descriptions, covering multiple modalities and tasks (He et al. 2024a). MedDr+RAD augments MedDr at inference with Retrieval-Augmented Diagnosis (RAD): for a test image, similar cases are retrieved and summarized as contextual guidance in the prompt (He et al. 2024b). GSCo (Generalist–Specialist Collaboration) is a two-stage framework that (i) builds a generalist GFM (MedDr) and lightweight specialist models, and (ii) performs collaborative inference via Mixture-of-Expert Diagnosis (MoED; using specialist predictions as references) and RAD (using specialists to retrieve similar cases) (He et al. 2024b). GSCo evaluates across a large multi-dataset benchmark; we include GSCo as a strong InternVL-based baseline along with its MedDr and MedDr+RAD components (He et al. 2024b; Chen et al. 2024).
Appendix D Additional Analysis
Performance with a State-of-the-Art LVLM Reranker
We evaluate whether a state-of-the-art large vision–language model (LVLM), used as an optional reranker, can improve performance on inconsistent-retrieval predictions. Specifically, we test the o3 reasoning model (OpenAI 2025) as a reranker: given a query image to classify and four retrieved image–caption pairs, o3 is asked to select the single pair that is most informative for predicting the correct label. We compare this reranking to four alternatives: (i) using the top-1 retrieved candidate (no reranking), (ii) selecting the model’s prediction with the highest confidence (maximum logit), (iii) using o3 with captions only (no images), and (iv) the Jomed aggregation strategy.
Our results show that o3 generally outperforms the maximum-logit baseline but remains inferior to Jomed’s logit-fusion aggregation. We evaluated performance on the Breast, Derma, and Retina datasets. For the remaining datasets—VinDr-PCXR, BRSET, and the visual question answering (VQA) datasets—we did not run this evaluation because the images and/or retrieved candidates come from restricted-access sources (e.g., MIMIC-CXR, VinDr-PCXR, and BRSET). See results in Table 8.
Jomed variants with medical pre-trained backbones
Although our work focuses on general-purpose LVLMs, we also evaluated a medical-specific baseline. We conducted the evaluation on four benchmarks—Breast, Derma, Retina, and VQA-RAD (closed-question subset). For the medical pre-trained baseline, we used Med-Flamingo and paired it with a medically pre-trained retriever (BiomedCLIP), using only one retrieval head (image retrieval without text retrieval). The performance gap between Med-Flamingo and our reader approach was small—0.01 in accuracy and 0.03 in F1. We hypothesize that this narrow gap may be due to Med-Flamingo’s relatively older LLaMA backbone (Touvron et al. 2023). Results are presented in Table 10. We also tested our general-purpose backbone with a medically pre-trained retriever, which showed improvements over FT RAG, with a similar pattern to the general-purpose retriever. Full results are presented in Table 9.
Retrieval Robustness Analyses.
We strive for our model to be robust in cases where the retrieval is noisy. Meaning, that the intrinsic ability in the model parameters is not overly dependent on the retrieval content. To test this property without ground truth retrieval labels, we evaluate the model performance with random candidates instead of searching according to cosine similarity. Overall as shown in Table 7 in Appendix D, we observed that for all datasets the model performance is above the reader-only, showing the model succeeded more than a model trained without any retrieval.
We also test the ability of the model to still function without any retrieved context. We do this to check whether some intrinsic capability is still maintained in the model’s reader without any retrieved information. This could also be helpful in cases where model users would like it to be versatile and support both retrieval-augmented prediction and non-retrieval-augmented prediction (e.g., in cases where no retrieved contexts were discovered). As shown in Table 7 in Appendix D, removing retrieval from Jomed resulted in performance comparable or slightly lower than the reader-only model.
Finally, we tested the model without the input image, relying only on the (image, report / caption) retrievals, to further assess the information held in retrieved data for predictions. In this scenario, we expected a substantial performance drop because the model no longer had direct knowledge of the patient’s condition, aside from what was provided by retrieved candidates. As a baseline, we used a simple random guess (coin flip), representing a scenario in which retrieval was not used effectively. As shown in Table 7 in Appendix D, our model’s performance exceeded this baseline by a large margin, with improvements of 0.32 in Accuracy, 0.28 in F1. These results indicate that the model learns from the retrieved data, yet it does not become overly dependent on it as we previously showed.
Appendix E Prompt Design
Here, we detail the prompts used in our experiments and analysis. To explore the utility of o3 as re-rankers of the most predictive (image, caption) pair, as described in Section D in Appendix D, we provide both models with a detailed prompt for the task, shown in Prompt 1.
Additionally, we present the prompts used to train Jomed for each of the datasets: Prompts 2, 4, 3, 6, 5, and the prompt for visual question answering tasks Prompt 7.
1 Task:
2
3 You are an expert assistant selecting the most informative reference for medical image classification.
4
5 Description:
6 A patients [MRI/CT/X-ray] image needs to be classified into one of the following categories: [List of possible labels].
7 You are given four candidate reference pairs, each consisting of a caption and its associated image, drawn from PubMed literature.
8 Your goal is to determine which single candidate provides the most useful information to help an AI model correctly classify the provided image.
9
10 Input Information:
11 - Medical image: The patients MRI/CT/X-ray image is provided here.
12 - Classification task: Classify the image into one of the following categories: [List of possible labels].
13 - Candidates:
14 - Candidate 0:
15 - Caption: [Caption 0]
16 - Image: [Image 0]
17 - Candidate 1:
18 - Caption: [Caption 1]
19 - Image: [Image 1]
20 - Candidate 2:
21 - Caption: [Caption 2]
22 - Image: [Image 2]
23 - Candidate 3:
24 - Caption: [Caption 3]
25 - Image: [Image 3]
26
27 Instructions:
28 1. For each candidate, carefully assess how informative its caption and image are for the current classification task.
29 2. Specifically, evaluate whether the candidate describes imaging features, findings, or clinical context relevant to distinguishing among the listed categories.
30 3. Compare all candidates and select the one that gives the clearest, most discriminative information to aid the classification.
31 4. Output only the number of the selected candidate in the following format:
32 Caption Number: [your selected number]
Listing 1: Prompt used for o3
[…]
Listing 2: Prompt used by Jomed for Breast classification
[…]
Listing 3: Prompt used by Jomed for Retina classification
[…]
Listing 4: Prompt used by Jomed for Derma classification
[…]
Listing 5: Prompt used by Jomed for Vindr-PCXR classification
[…]
Listing 6: Prompt used by Jomed for BRSET classification
[…]
Listing 7: Prompt used by Jomed for visual question answering tasks

### v3

@@
-AAAI Press Formatting Instructions for Authors Using — A Guide
+Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis
 Abstract
-Clinical decision-making often involves interpreting images (e.g., radiology) for making diagnoses.
-Retrieving relevant visual information from medical literature and hospital records could enhance diagnostic accuracy.
-In this paper, we develop a model in which a multimodal retriever is jointly optimized with an LVLM for medical diagnosis, unlike standard RAG where LVLM error signal is not propagated down to the retriever.
-We show that using only general-purpose backbones, with only lightweight fine-tuning, our model is able to achieve competitive results with medically-pretrained models across clinical multi-label classification and visual question answering tasks.
-In a novel analysis, we additionally find that in many cases different top retrieved images each lead to different predictions for a given target, and that these cases are empirically challenging for all models, even for non-retrieval models.
-Our joint retrieval optimization significantly improves these challenging cases over standard RAG.
-However, oracle analysis reveals that while the correct diagnosis is frequently achievable using one of the top retrieved images, in practice there is a large performance gap from the oracle, and rerankers using frontier LVLMs do not close this gap–leaving ample room for improvement by future methods.Code available at: https://github.com/Nirmaz/JOMED.
-Introduction
-Figure 1: We jointly optimize a multimodal retriever and a Large Vision-Language Model (LVLM) for medical tasks.
-Our model achieves competitive results without resource-intensive medical pre-training and significantly improves performance on challenging cases where different retrieved images lead to inconsistent retrieval prediction.
+Retrieving relevant visual and textual information from medical literature and hospital records can enhance diagnostic accuracy for clinical image interpretation.
+We develop a multimodal retrieval model jointly optimized with an LVLM for medical diagnosis, unlike standard RAG which doesn’t backpropagate LVLM errors to the retriever.
+Using only general-purpose backbones with lightweight fine-tuning, our model achieves competitive results with medically-pretrained models on clinical classification and VQA tasks.
+In a novel analysis, we find that different top-retrieved images often yield different predictions for the same target, and that these cases are challenging for all models, even for non-retrieval models.
+Our joint retrieval optimization significantly improves these cases over standard RAG.
+However, oracle analysis reveals that while the correct diagnosis is frequently achievable using one of the top retrieved images, in practice there is a large performance gap from the oracle, and rerankers using frontier LVLMs do not close this gap—leaving ample room for improvement by future methods.
+Code available at https://github.com/Nirmaz/JOMED.
+Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis
+Nir Mazor1, Tom Hope1,2 1School of Computer Science and Engineering, The Hebrew University of Jerusalem 2The Allen Institute for AI (AI2)
+1 Introduction
+Figure 1: We jointly optimize a multimodal retriever and a Large Vision-Language Model for medical tasks.
+We achieve competitive results without resource-intensive medical pre-training and significantly improve performance on challenging cases where different retrieved images lead to inconsistent retrieval prediction.
 Inferring diagnoses from medical imagery is a fundamental part of clinical decision-making.
-Large Vision Language Models (LVLMs) have been widely explored for medical diagnosis (Thawakar et al.
-2024; Wu et al.
-2023; Zhang et al.
-2023b; Moor et al.
-2023; Li et al.
-2023).
-To improve the performance of LVLMs in the medical field, retrieval augmentation has been adopted and has shown promising results, providing more accurate and also explainable methods (He et al.
-2024a; Xia et al.
-2024b, a; Wu et al.
-2025).
-In this work, we explore a lightweight fine-tuning approach for taking a general-purpose LVLM and optimizing it jointly with a general-purpose multimodal retriever.
-Different from the common retrieval augmented generation (RAG) setup, we perform joint optimization of the LVLM and retriever.
-This joint optimization teaches the LVLM to utilize retrieved knowledge for making accurate medical predictions while at the same time tuning the multimodal retriever to retrieve information that leads to correct predictions by the LVLM.
-The retriever model retrieves images as well as captions and clinical reports from both literature and hospital records.
-Our method involves sequential multimodal training with a dual-head retriever architecture and a retrieval loss that enhances the retriever’s ability to provide more predictive context to the LVLM.
-Poorly optimized retrieval mechanisms can mislead models (Yoran et al.
-2023; Sun et al.
-2024).
-Figure 2 demonstrates this problem with an ultrasound image that shows no breast cancer.
-When the retrieval system selects an irrelevant reference image showing cancer features, the model incorrectly identifies cancer in the input image.
-When we jointly optimize the retriever to find more relevant images and text, the model’s prediction flips to the correct diagnosis.
-Using our approach, we are able to achieve competitive results in medical image classification and visual question answering (VQA).
+Large Vision Language Models (LVLMs) have been widely explored for medical diagnosis Thawakar et al.
+(2024); Wu et al.
+(2023); Zhang et al.
+(2023b); Moor et al.
+(2023); Li et al.
+(2023).
+To improve the performance of LVLMs in the medical field, retrieval augmentation (RAG) has been adopted and has shown promising results, providing more accurate and also explainable methods He et al.
+(2024a); Xia et al.
+(2024b, a); Wu et al.
+(2025).
+In this work, we explore a lightweight fine-tuning approach for taking a general-purpose LVLM and optimizing it together with a general-purpose multimodal retriever for medical diagnosis tasks.
+Different from the common multimodal RAG setup, we perform joint optimization of the LVLM and retriever.
+This teaches the LVLM to utilize retrieved knowledge—images, captions and clinical reports from both literature and hospital records—while at the same time tuning the retriever to find information that leads to correct predictions by the LVLM.
+Poorly optimized retrieval mechanisms can mislead models Yoran et al.
+(2023); Sun et al.
+(2024); as shown in Figure 2, a RAG model incorrectly diagnoses a benign ultrasound when using a cancerous retrieved image, but joint optimization leads to correct retrieval and prediction.
+Our method involves sequential multimodal training with a dual-head retriever architecture and a customized retrieval loss and visual question answering (VQA) training recipe.
+Our method achieves competitive results in medical image classification and VQA.
 Importantly, our focus in this work is not to chase SOTA results but to shed light on several interesting observations.
-One, that a simple and effective joint LVLM-retriever optimization mechanism can lead to a substantial boost in results over standard multimodal RAG and should thus be more widely adopted for medical diagnosis.
-Second, that this mechanism achieves competitive results by using models with no medical pre-training (for neither the LVLM nor the retriever) and only lightweight fine-tuning; this resonates with recent findings showing general-purpose LVLMs can rival their medical counterparts (Jeong et al.
-2024) and has potentially important implications considering the resource-intensive nature of pre-training processes.
-Third, previous work on joint training in the general domain required large-scale pre-training followed by few-shot task-specific adaptation (Izacard et al.
-2023; Hu et al.
-2023; Lin et al.
-2024).
-In contrast, our novelty is in conducting joint training directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples).
-This provides first evidence about the feasibility and utility of lightweight data-efficient joint training directly on downstream tasks as opposed to in the pre-training setting.
-We are also the first to explore and show the utility of joint training in the multimodal medical domain.
+One, that a simple and effective joint LVLM-retriever optimization mechanism can lead to a substantial boost in results over current multimodal RAG methods and should thus be more widely adopted for medical diagnosis.
+Second, that this mechanism achieves competitive results by using models with no medical pre-training (for neither the LVLM nor the retriever) and only lightweight fine-tuning.
+This resonates with recent findings showing general-purpose LVLMs can rival their medical counterparts Jeong et al.
+(2024) and has potentially important implications considering the resource-intensive nature of pre-training processes.
+Third, previous work on joint optimization in the general domain required large-scale pre-training followed by few-shot task-specific adaptation Izacard et al.
+(2023); Hu et al.
+(2023); Lin et al.
+(2024).
+In contrast, we conduct joint optimization directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples).
+This provides first evidence about the feasibility and utility of lightweight data-efficient joint optimization directly on downstream tasks as opposed to in the pre-training setting.
+We are also the first to explore and show the utility of joint optimization in the multimodal medical domain.
 Fourth and importantly, we conduct a novel analysis that finds that our method helps in particular to address a class of errors we term inconsistent retrieval predictions.
@@
 These cases commonly occur across our experiments and are substantially more difficult for models, both retrieval-augmented and non-retrieval models.
-This instability with respect to different retrieval candidates leads to high prediction entropy/uncertainty across classes, degrading not only overall RAG results but also their reliability (Lambert et al.
-2022).
-As we show, our joint training mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases.
+This instability with respect to different retrieval candidates leads to high prediction entropy/uncertainty across classes, degrading not only overall RAG results but also their reliability Lambert et al.
+(2022).
+As we show, our joint optimization mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases.
 We further discover that for a large proportion of these inconsistent cases, at least one retrieved candidate enables the model to predict the correct answer in an oracle setting in which we provide the model with the correct retrieved candidate.
 This suggests useful information is being retrieved but is often lost among less helpful candidates.
-To address this, we explored using the powerful multimodal model o3 (OpenAI 2025) as a reranker to identify the most predictive instances.
+To address this, we explored using the powerful multimodal model o3 OpenAI (2025) as a reranker to identify the most predictive instances.
 We found it did not close the performance gap to the oracle and was comparable or inferior to our method, leaving significant room for future improvement.
 Our contributions are three-fold:
-  • We develop a method where a multimodal retriever and a Large Vision-Language Model (LVLM) are jointly optimized for medical classification and VQA.
-Unlike standard Retrieval-Augmented Generation (RAG), our method propagates error signal from the LVLM down to the retriever while training on downstream tasks.
-  • We show that our method achieves competitive results compared to medically pre-trained models by using only general-purpose backbones.
-Unlike previous general-domain joint training methods that require pre-training, we conduct joint training directly on downstream tasks with only lightweight data-efficient fine-tuning.
-  • In a novel analysis we identify inconsistent retrieval predictions, challenging cases where different retrieved images lead to different predictions for the same query.
+  • We jointly optimize a multimodal retriever and an LVLM for medical classification and VQA.
+We use only general-purpose backbones without any medical pretraining, and show that by retrieving medical information and co-optimizing the retriever and LVLM we can close much of the gap to expensive medically pre-trained models.
+  • Unlike previous general-domain joint optimization that required pre-training, we conduct joint optimization directly on downstream tasks with only lightweight data-efficient fine-tuning.
+  • We identify inconsistent retrieval predictions, challenging cases where different retrieved images lead to different predictions for the same query.
 Our joint optimization method significantly improves performance on these cases when compared to standard RAG.
-Figure 2: Joint training impacts inconsistent retrieval predictions.
-After joint training, retrieved images show greater alignment with query image labels.
-In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after joint training, the retrieved image is less directly related to the wrong label (c).
-For Retina, joint training changed retrieval from a different modality (e) to be more similar to the query (d).
-In VQA-RAD, retrieval shifted from an unrelated medical condition (h) to an image depicting the condition of the query image.
-Related Work
-Retrieval Augmented Joint Training has been explored primarily in unimodal (text only) NLP with work such as ATLAS (Izacard et al.
-2023), which conducted large-scale pretraining of both reader and retriever models and evaluated in zero-shot and few-shot scenarios in the general domain.
-REVEAL (Hu et al.
-2023) extended this to the multimodal setting, achieving state-of-the-art results in visual tasks in the general domain with extensive pre-training.
-(Shi et al.
-2023) proposed training textual retrievers based on reader performance, influencing subsequent joint-training methods (Lin et al.
-2024).
-introduced another joint training variant and proposed sequential training: first training the reader with retrieval augmentation, then training the retriever.
+Figure 2: joint optimization impacts inconsistent retrieval predictions.
+After joint optimization, retrieved images show greater alignment with query image labels.
+In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after joint optimization, the retrieved image is less directly related to the wrong label (c).
+In VQA-RAD, retrieval shifted from an unrelated medical condition (e) to an image depicting the condition of the query image.
+2 Related Work
+Retrieval Augmented joint optimization has been explored primarily in unimodal (text only) NLP with work such as ATLAS Izacard et al.
+(2023), which conducted large-scale pretraining of both reader and retriever models and evaluated in zero-shot and few-shot scenarios in the general domain.
+REVEAL Hu et al.
+(2023) extended this to the multimodal setting, achieving state-of-the-art results in visual tasks in the general domain with extensive pre-training.
+Shi et al.
+(2023) proposed training textual retrievers based on reader performance, influencing subsequent joint-optimization methods.
+Lin et al.
+(2024) introduced another joint optimization variant and proposed sequential training: first training the reader with retrieval augmentation, then training the retriever.
 Siriwardhana et al.
-(2023) demonstrated joint training’s effectiveness for domain adaptation.
-Our work differs by being the first to explore joint training optimized directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training.
-Our extensive novel analyses are the first to shed light on inconsistent predictions and how joint training on downstream tasks helps address them.
+(2023) demonstrated joint optimization’s effectiveness for domain adaptation.
+Our work differs by being the first to explore joint optimization optimized directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training.
+Our extensive novel analyses are the first to shed light on inconsistent predictions and how joint optimization on downstream tasks helps address them.
 Multimodal Retrieval Augmentation in Medical Applications.
-Multimodal retrieval augmentation began with encoder-based architectures (Yuan et al.
-2023), which integrated retrieved text and images with query modalities.
-With Large Vision-Language Models (LVLMs), (He et al.
-2024a) proposed retrieving relevant labeled examples during inference.
-(Xia et al.
-2024b) addressed LVLMs’ misalignment with medical knowledge using retrieval-augmented generation to fetch similar medical reports, employing contrastive learning and context selection strategies.
-(Xia et al.
-2024a) proposed domain-aware retrieval for diverse medical data sources, while (Wu et al.
-2025) combined retrieval augmentation with knowledge graphs to reduce hallucination.
-Method
-Figure 3: Jomed combines a general-purpose VLLM with a general-purpose multi-modal retriever in a two-phase approach.
-First, with a frozen retriever, we augment training by prepending retrieved (image, caption/report) pairs to original queries, training the LVLM on these augmented prompts.
-Second, with the LVLM frozen, we sequentially train the dual-head multi-modal retriever using Sharp Perplexity Distillation loss (KL-divergence between the distribution derived from the model logits and the distribution derived from the retrieved candidates’ similarity).
+Multimodal retrieval augmentation began with encoder-based architectures Yuan et al.
+(2023), which integrated retrieved text and images with query modalities.
+With Large Vision-Language Models (LVLMs), He et al.
+(2024a) proposed retrieving relevant labeled examples during inference.
+Xia et al.
+(2024b) addressed LVLMs’ misalignment with medical knowledge using retrieval-augmented generation to fetch similar medical reports, employing contrastive learning and context selection strategies.
+Xia et al.
+(2024a) proposed domain-aware retrieval for diverse medical data sources, while Wu et al.
+(2025) combined retrieval augmentation with knowledge graphs.
+3 Method
+Figure 3: The two-phase Jomed training.
+The LVLM is first trained with a frozen retriever on augmented prompts containing retrieved image–caption/report pairs.
+With the LVLM frozen, the dual-head multimodal retriever is optimized using a KL divergence computed over a selected subset of the model’s logits.
+An LLM converts open-ended questions into closed form (during training only), which improves results.
 Overview.
-In this section, we present our methodology, named Jomed (Joint Optimization of Multimodal Retrieval for Enhanced Medical Diagnosis).
-Our task involves classifying medical images and medical visual question answering.
-Jomed consists of two main components: a multimodal retriever and a reader.
-We are given an input patient image and a diagnostic question.
-The multimodal retriever is tasked with identifying relevant medical knowledge from either literature or hospital records—in the form of images and associated captions or hospital reports that provide predictive information.
-The reader then analyzes the retrieved candidates along with the diagnostic question and generates an answer.
-We jointly optimize the reader to better utilize retrieved candidates, and the multimodal retriever to improve its ability to retrieve informative candidates for the reader.
-Figure 3 illustrates our method.
-The multimodal retriever has two retrieval heads, textual and visual, for finding candidates in an index of image-caption pairs derived from literature and hospital records.
-Given an image query, these retrievers find candidates that are then passed to the reader.
-Unlike previous work (He et al.
-2024a; Xia et al.
-2024b, a) our method does not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip (Xiao, Mastrapas, and Wang 2024).
+In this section, we present our methodology, termed Jomed (Joint Optimization of Multimodal Retrieval for Enhanced Medical Diagnosis).
+Our task involves medical image classification and visual question answering.
+Jomed comprises two main components: a multimodal retriever and a reader.
+Given an input patient image and a diagnostic question, the multimodal retriever identifies relevant medical knowledge—images and their associated captions or hospital reports—that provide predictive information.
+The reader then analyzes the retrieved candidates along with the question to generate an answer.
+We jointly optimize the reader to better leverage the retrieved information and the multimodal retriever to enhance its ability to select informative candidates for the reader.
+Figure 3 illustrates our framework.
+Unlike previous work He et al.
+(2024a); Xia et al.
+(2024b, a) our method does not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip Xiao et al.
+(2024).
 The selected LVLMs can process multiple images and texts, allowing us to incorporate both the retrieved image and the retrieved caption/report into the model context.
-We also compare our results to Med-Flamingo (Moor et al.
-2023).
-MedFlamingo is a medical LVLM which is also able to process multiple images, unlike other medical LVLM available at the time of conducting our experiments.
-Finally, we also explore the utility of o3 (OpenAI 2025) as a reranker.
+We also compare our results to Med-Flamingo Moor et al.
+(2023).
+Med-Flamingo is a medical LVLM which is also able to process multiple images, unlike other medical LVLM available at the time of conducting our experiments.
 We now elaborate on the details.
-Reader Retrieval Augmentation Fine-Tuning
+3.1 Reader Retrieval Augmentation Fine-Tuning
 We first fine-tune the reader with retrieval augmentation.
 This step aims to achieve two main goals: improving the reader’s performance on the dataset and teaching the reader to effectively utilize retrieved (image, text) pairs in context.
-For an image, question tuple $(i_{d},q_{d})$ $d\in\{1,\ldots,D\}$ from dataset $D$ , we retrieve $K$ relevant tuples of images and texts (reports or captions) $(i_{k},t_{k})$ $k\in\{1,\ldots,K\}$ using our multimodal retriever module.
-These retrieved pairs are prepended to the image-question pair $(i_{d},q_{d})$ to create augmented inputs.
-The augmented input consists of the retrieved context followed by the target question-image pair: $(i_{k},t_{k}),(i_{d},q_{d})$ $k\in\{1,\ldots,K\}$.
-To the best of our knowledge, we are the first to use both the retrieved image and caption/report in context within medical RAG for LVLMs.
+For an image-question tuple $(i_{d},q_{d})$ where $d\in\{1,\ldots,D\}$ from dataset $D$, we retrieve $K$ relevant tuples of images and texts (reports or captions) $(i_{k},t_{k})$ where $k\in\{1,\ldots,K\}$ using our multimodal retriever module.
+To simplify notation, we denote $z_{k}=(i_{k},t_{k})$ as the $k$-th retrieved image-text pair and $q_{d}=(i_{d},q_{d})$ as the query pair for example $d$.
+These retrieved pairs are prepended to the query pair to create augmented inputs.
+The augmented input consists of the retrieved context followed by the target question-image pair: $z_{k}\circ q_{d}$ for $k\in\{1,\ldots,K\}$.
 We use supervised fine-tuning on the augmented input, minimizing the negative log-likelihood to predict the answer $a_{d}$:
-$\mathcal{L}(\theta)=-\sum_{d=1}^{D}\sum_{k=1}^{K}\log p_{\theta}(a_{d}|(i_{k},t_{k})\circ(i_{d},q_{d}))$
-where $\theta$ represents the parameters of the LVLM, and $\circ$ denotes the concatenation operation in the context window.
-Sequential Multimodal Retriever Fine-Tuning
+$\mathcal{L}(\theta)=-\sum_{d=1}^{D}\sum_{k=1}^{K}\log p_{\theta}(a_{d}\mid z_{k}\circ q_{d})$
+where $\theta$ represents the parameters of the LVLM, and $\circ$ denotes the concatenation operation.
+3.2 Sequential Multimodal Retriever Fine-Tuning
 Our dual-head multimodal retriever has a text-based head for retrieving relevant (image, text) pairs based on textual similarity to the query, and an image-based head for retrieving pairs based on visual similarity.
-To train this multimodal retriever, we measure how each retrieved (image, text) pair contributes to the LVLM’s performance, optimizing the retriever to prioritize high-contribution candidates.
-We adopt a sequential training strategy, first optimizing the text retrieval head followed by the image retrieval head.
+To train this retriever, we adopt a sequential training strategy, first optimizing the text retrieval head followed by the image retrieval head.
 During this process, the reader model remains frozen, focusing solely on improving the retriever’s embedding space.
-We formulate a Sharp Perplexity Loss as a simple variant of the perplexity distillation loss (Izacard et al.
-2023).
 Similarly to Izacard et al.
-(2023) we minimize the KL-divergence between the LVLM’s posterior distribution over retrieved pairs and the retriever’s distribution, however we do so directly on downstream tasks with only lightweight fine-tuning in the data-efficient regime without pre-training.
-The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved (image, text) pair $(\mathbf{i}_{k},\mathbf{t}_{k})$ and a query image-question pair $(\mathbf{i},\mathbf{q})$:
-$p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q})),$
-where $p_{\text{LVLM}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q}))$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
+(2023) we minimize the KL divergence between the LVLM’s posterior distribution over retrieved pairs and the retriever’s distribution, however we do so directly on downstream tasks with only lightweight fine-tuning in the data-efficient regime without pre-training, and we customize the loss and the training recipe for open-ended questions leading to empirical gains.
+Let $z_{k}=(\mathbf{i}_{k},\mathbf{t}_{k})$ denote the $k$-th retrieved image-text pair and $q=(\mathbf{i},\mathbf{q})$ denote the query pair.
+The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved pair $z_{k}$ and query $q$:
+$p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q),$
+where $p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q)$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
 We sharpen the LVLM’s output distribution by restricting it to only the relevant class tokens from the benchmark’s possible answers.
-Specifically, we extract the model’s logits, select only the relevant tokens for each class, and apply a softmax to obtain a more discriminative distribution (for VQA tasks, we use yes/no tokens and skip open questions.).
+Specifically, we extract the model’s logits, select only the relevant tokens for each class, and apply a softmax to obtain a more discriminative distribution.
 We denote this class-restricted distribution as $p_{\text{LVLM}_{C}}$ where $C=\{c_{1},c_{2},c_{3},\ldots\}$ is the set of chosen class tokens.
 The normalized posterior $p_{k}$ is formulated as:
-$p_{k}=\frac{\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ(\mathbf{i},\mathbf{q})))}{\sum_{i=1}^{K}\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid(\mathbf{i}_{i},\mathbf{t}_{i})\circ(\mathbf{i},\mathbf{q})))},$
+$p_{k}=\frac{\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid z_{k}\circ q))}{\sum_{i=1}^{K}\exp(\log p_{\text{LVLM}_{C}}(\mathbf{a}\mid z_{i}\circ q))},$
 where $K$ is the total number of retrieved pairs.
 The retriever’s distribution is defined by:
-$p_{\text{RETR}}((\mathbf{i},\mathbf{t})\mid(\mathbf{i},\mathbf{q}))=\frac{\exp(s((\mathbf{i},\mathbf{t}),(\mathbf{i},\mathbf{q}))/\tau)}{\sum_{k=1}^{K}\exp(s((\mathbf{i}_{k},\mathbf{t}_{k}),(\mathbf{i},\mathbf{q}))/\tau)},$
-where $s((\mathbf{i},\mathbf{t}),(\mathbf{i},\mathbf{q}))$ is the similarity score between the pair and the query, and temperature $\tau$ controls distribution sharpness.
+$p_{\text{RETR}}(z\mid q)=\frac{\exp(s(z,q)/\tau)}{\sum_{k=1}^{K}\exp(s(z_{k},q)/\tau)},$
+where $s(z,q)$ is the similarity score between the pair and the query, and temperature $\tau$ controls distribution sharpness.
 For the text retrieval head, the score is computed using the dot product between the index caption/report embeddings and the query image embedding, while for the image retrieval head, we use the index image embeddings.
-Finally, for our loss we compute the KL-divergence between retriever and reader distributions $\text{KL}(p_{\text{LVLM}_{C}}\|p_{\text{RETR}})$:
-$\sum_{k=1}^{K}p_{\text{LVLM}_{C}}((\mathbf{i}_{k},\mathbf{t}_{k}))\log\left(\frac{p_{\text{LVLM}_{C}}((\mathbf{i}_{k},\mathbf{t}_{k}))}{p_{\text{RETR}}((\mathbf{i}_{k},\mathbf{t}_{k}))}\right)$
-Inference
+For our loss we compute the KL-divergence between retriever and reader $\text{KL}(p_{\text{LVLM}_{C}}\|p_{\text{RETR}})$:
+$\sum_{k=1}^{K}p_{\text{LVLM}_{C}}(z_{k})\log\left(\frac{p_{\text{LVLM}_{C}}(z_{k})}{p_{\text{RETR}}(z_{k})}\right)$
+Finally, for visual question answering (VQA), we use the o3 model to convert open-ended questions into closed-ended ones (yes/no).
+See prompt in Appendix E and analysis of approach contribution in Appendix D.1.
+We apply this only during training [^1] Run once per question; the total cost is only a few dollars., without modifying questions at inference time.
+Reformatted questions will be made publicly available.
+This reformulation improves results, akin to restricting the LVLM output distribution.
+3.3 Inference
 Following Shi et al.
-(2023) for a given (image, question) pair, we retrieve the top-$k$ relevant (image, text) chunks using the image as the query.
+(2023) for a given query pair $q=(\mathbf{i},\mathbf{q})$, we retrieve the top-$K$ relevant chunks $z_{k}=(\mathbf{i}_{k},\mathbf{t}_{k})$ using the image as the query.
 Each chunk is prepended to the question, and the LVLM computes predictions for the augmented prompts in parallel.
 The final output probability is:
-$p_{\text{LVLM}}(a\mid\mathbf{q})=\sum_{k=1}^{K}p_{\text{LVLM}}(a\mid(\mathbf{i}_{k},\mathbf{t}_{k})\circ\mathbf{q})\cdot p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q}),$
-where $p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q})$ is:
-$p_{\text{R}}((\mathbf{i}_{k},\mathbf{t}_{k})\mid\mathbf{q})=\frac{\exp(s(\mathbf{q},(\mathbf{i}_{k},\mathbf{t}_{k})))}{\sum_{j=1}^{K}\exp(s(\mathbf{q},(\mathbf{i}_{j},\mathbf{t}_{j})))},$
-and $s(\mathbf{q},(\mathbf{i}_{k},\mathbf{t}_{k}))$ is the similarity score between the query and the retrieved chunk.
-Backbone Model Breast Derma Retina VinDr-PCXR BRSET Mean
+$p_{\text{LVLM}}(a\mid q)=\sum_{k=1}^{K}p_{\text{LVLM}}(a\mid z_{k}\circ q)\cdot p_{\text{R}}(z_{k}\mid q),$
+where $p_{\text{R}}(z_{k}\mid q)$ is:
+$p_{\text{R}}(z_{k}\mid q)=\frac{\exp(s(q,z_{k}))}{\sum_{j=1}^{K}\exp(s(q,z_{j}))},$
+and $s(q,z_{k})$ is the similarity score between the query and the retrieved chunk.
+Backbone (Size) Model Breast Derma Retina VinDr-PCXR BRSET Mean
 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
-Qwen2-vl Reader .83 .77 .68 .27 .60 .44 .48 .08 .34 .23 .59 .36
+LLaVA (7B) MMed-RAG .85 .84 .75 .30 .63 .46 .55 .11 .42 .30 .64 .40
+Qwen2-vl (7B) Reader .83 .77 .68 .27 .60 .44 .48 .08 .34 .23 .59 .36
+RAD .84 .79 .43 .34 .55 .40 .57 .09 .40 .25 .56 .37
 FT RAG .85 .82 .71 .42 .62 .48 .55 .09 .48 .27 .64 .42
+Jomed + PDist .86 .83 .73 .47 .63 .49 .57 .09 .47 .29 .65 .43
 Jomed .87 .84 .76 .50 .65 .50 .57 .14 .49 .37 .67 .47
 Jomedoracle .87 .84 .85 .64 .69 .51 .64 .14 .52 .41 .71 .51
-Pixtral Reader .82 .77 .75 .52 .55 .44 .48 .08 .44 .33 .61 .43
+Pixtral (12B) Reader .82 .77 .75 .52 .55 .44 .48 .08 .44 .33 .61 .43
+RAD .85 .80 .75 .47 .56 .41 .48 .09 .47 .30 .62 .41
 FT RAG .88 .85 .79 .60 .57 .47 .49 .09 .47 .33 .64 .47
+Jomed + PDist .88 .84 .80 .62 .60 .47 .50 .09 .45 .35 .65 .47
 Jomed .90 .87 .80 .62 .60 .51 .56 .14 .51 .37 .67 .50
 Jomedoracle .93 .90 .83 .67 .63 .54 .67 .15 .57 .41 .73 .53
-Backbone Model VQA-RAD SLAKE PathVQA Mean
+Backbone (Size) Model VQA-RAD SLAKE PathVQA Mean
 Closed Open Closed Open Closed Open Closed Open
-Qwen2-vl Reader .73 .41 .84 .80 .87 .25 .81 .49
+LLaVA (7B) MMed-RAG .74 .39 .87 .81 .90 .31 .84 .50
+Qwen2-vl (7B) Reader .73 .41 .84 .80 .87 .25 .81 .49
 FT RAG .76 .45 .88 .81 .91 .33 .85 .53
-Jomed .78 .47 .90 .83 .93 .37 .87 .56
+Jomed + PDist .77 .45 .89 .81 .92 .32 .86 .53
+Jomed .79 .48 .90 .84 .93 .38 .87 .57
 Jomedoracle .86 - .92 - .95 - .91 -
-Pixtral Reader .72 .38 .83 .80 .87 .26 .81 .48
+Pixtral (12B) Reader .72 .38 .83 .80 .87 .26 .81 .48
 FT RAG .74 .41 .88 .81 .88 .31 .83 .51
-Jomed 76 .43 .90 .82 .90 .35 .85 .53
+Jomed + PDist .75 .41 .89 .81 .89 .32 .84 .51
+Jomed .76 .45 .90 .84 .90 .36 .85 .55
 Jomedoracle .88 - .92 - .95 - .92 -
-Table 1: Results for classification (a) and visual question answering (b).
-Joint training consistently leads to the best results, with up to a 10-point increase in F1 score compared to fine-tuned (FT) RAG for multi-label classification.
-For visual question answering, we evaluate closed questions using the exact match metric and open questions using token recall.
-Using an oracle reranker demonstrates that even larger improvements are achievable with Jomed’s top-retrieved images in some datasets.
-Note that in the open question scenario, we do not evaluate the oracle because there is no clear discriminative boundary between different answers to determine the correct one, as there is in classification and closed question answering tasks.
-Using an oracle reranker shows that larger improvement is achievable with Jomed’s top retrieved images in some datasets.
-Experiments
-Medical
-Pretrained
+Table 1: Results for classification (a) and VQA (b) with non-medically pretrained LVLM backbones (see Tables 3, 4 for comparisons to with medical pre-training; MMed-RAG’s retrieval component is medically pre-trained).
+joint optimization consistently leads to the best results.
+For VQA, we evaluate closed questions using the exact match metric and open questions using token recall.
+Using an oracle reranker (for classification and closed VQA; §4.2) demonstrates that even larger improvements are achievable with Jomed’s top-retrieved images in some datasets.
+MMed-RAG
+LLaVaMed
+variants
+MedDr
+variants
+GSCo
+MedVInT
+variants
+Medical Pre-Training
 Size
-Model Breast Derma
-177K MedVInT-TE .88 .78
-177K MedVInT-TD .90 .80
-177K MedDrInternVL .72 -
-177K MedDr + RADInternVL .88 -
-255K GSCoInternVL .93 -
-0 Jomed Qwen2-vl .87 .76
-0 JomedPixtral .90 .80
-Table 2: Accuracy of Jomed and prior reported results of medical pre-trained models for multiclass classification.
-Medical
-Pretrained
-Size
-Model VinDr-PCXR BRSET
-255K MedDrInternVL .08 .08
-255K GSCoInternVL .09 .33
-0 Jomed Qwen2-vl .14 .37
-0 JomedPixtral .14 .37
-Table 3: F1 Comparison of Jomed to prior reported results of medical pre-trained models for multi-label classification.
-Medical
-Pretrained
-Size
+600K 600K 255K 255K 177K
+Table 2: Medical pre-training size for LVLMs under extensive medical pretraining.
+4 Experiments
+Model Breast Derma VinDr BRSET
+MMed-RAGLLaVA-Med .89 .79 .11 .33
+MedVInT-TE .88 .78 - -
+MedVInT-TD .90 .80 - -
+MedDrInternVL .72 - .08 .08
+MedDr + RADInternVL .88 - - -
+GSCoInternVL .93 - .09 .33
+Jomed Qwen2-vl .87 .76 .14 .37
+JomedPixtral .90 .80 .14 .37
+Table 3: Comparison to prior reported results of medical pre-trained models, for binary and multiclass classification (Breast and Derma; accuracy), and for multi-label classification (VinDr-PCXR and BRSET; F1).
 Model SLAKE PathVQA
-Closed Open Closed Open Open
-600K BiomedGPT-B .90 .85 .88 .28
-600K LLaVA-MedLLaVA .83 .85 .91 .38
-600K LLaVA-MedVicuna .85 .83 .92 .39
-600K LLaVA-MedBioMedCLIP .87 .87 .91 .39
-0 JomedQwen2-vl .90 .83 (.84) .93 .37 (.37)
-0 JomedPixtral .90 .82 (.82) .90 .35 (.35)
-Table 4: Comparison of Jomed to medical pre-trained models for visual question answering.
-We report the token recall metric, reported for LLaVA-Med variants and token F1 reported for BiomedGPT (in red).
-Experimental Setup
-Our datasets include real-world hospital datasets (BRSET (Nakayama et al.
-2024) and VinDr‑PCXR (Pham, Tran, and Nguyen 2022)) alongside a variety of classification and visual question answering datasets.
+Closed Open Closed Open
+MMed-RAGLLaVA-Med .89 .84 .92 .39
+BiomedGPT-B .90 .85 .88 .28
+LLaVA-MedLLaVA .83 .85 .91 .38
+LLaVA-MedVicuna .85 .83 .92 .39
+LLaVA-MedBioMedCLIP .87 .87 .91 .39
+JomedQwen2-vl .90 .84 (.84) .93 .38 (.37)
+JomedPixtral .90 .84 (.82) .90 .36 (.35)
+Table 4: Comparison to medical pre-trained models for visual question answering. We report the token recall metric, reported for LLaVA-Med variants and token F1 reported for BiomedGPT (in red).
+4.1 Experimental Setup
+Our datasets include real-world hospital datasets (BRSET Nakayama et al.
+(2024) and VinDr‑PCXR Pham et al.
+(2022)) alongside a variety of classification and visual question answering datasets.
 We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA).
 Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive.
-Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed (Casey 2022).
+Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed Casey (2022).
 Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study.
-Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge to compensate for sparse training samples.
+Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge.
 We consider three types of classification tasks: binary, multi-class, and multi-label.
-These include BreastMNIST (nickname: Breast; breast ultrasound imaging, binary) (Al-Dhabyani et al.
-2020), DermaMNIST (nickname: Derma; pigmented skin lesion images, multi-class) (Tschandl, Rosendahl, and Kittler 2018), RetinaMNIST (nickname: Retina; retinal fundus images with multi-label disease annotations) (Liu et al.
-2022), and two challenging real-world multi-label datasets: VinDr-PCXR (a chest X-ray dataset with 15 labels) (Pham, Tran, and Nguyen 2022) and BRSET (an ophthalmology dataset with 14 labels) (Nakayama et al.
-2024).
-For visual question answering, we use widely adopted benchmarks: VQA‑RAD (Lau et al.
-2018), SLAKE‑English (Liu et al.
-2021), and PathVQA (He et al.
-2020).
-Full details are provided in Appendix A.
-Retrieval augmentation is supported by an external index constructed from PubMed and medical records: PMC‑OA (Lin et al.
-2023), MIMIC‑CXR (Johnson et al.
-2019), and ROCO (Rückert et al.
-2024).
-Full descriptions are available in Appendix A.
-Our experiments leverage two backbone models—Pixtral (12B) (Agrawal et al.
-2024) and Qwen2‑vl (7B parameters) (Wang et al.
-2024)—with jina‑clip‑v1 (Xiao, Mastrapas, and Wang 2024) serving as the general-purpose retriever’s visual head, embedding both the texts and images of the retrieval index.
-All models were trained in a low-resource setting, with all runs performed on a single L40s GPU.
+These include BreastMNIST (“Breast”; breast ultrasound imaging, binary) Al-Dhabyani et al.
+(2020), DermaMNIST (“Derma”; pigmented skin lesion images, multi-class) Tschandl et al.
+(2018), RetinaMNIST (“Retina”; retinal fundus images with multi-class disease annotations) Liu et al.
+(2022), and two challenging real-world multi-label datasets: VinDr-PCXR (a chest X-ray dataset with 15 labels) Pham et al.
+(2022) and BRSET (an ophthalmology dataset with 14 labels) Nakayama et al.
+(2024).
+For visual question answering, we use widely adopted benchmarks: VQA‑RAD Lau et al.
+(2018), SLAKE‑English Liu et al.
+(2021), and PathVQA He et al.
+(2020).
+Full details are provided in Appendix A.1.
+Retrieval augmentation is supported by an external index constructed from PubMed and medical records: PMC‑OA Lin et al.
+(2023), MIMIC‑CXR Johnson et al.
+(2019a), and ROCO Rückert et al.
+(2024).
+Full descriptions are available in Appendix A.2.
+Our experiments leverage two backbone models—Pixtral (12B) Agrawal et al.
+(2024) and Qwen2‑vl (7B parameters) Wang et al.
+(2024)—with jina‑clip‑v1 Xiao et al.
+(2024) serving as the general-purpose retriever’s visual head, embedding both the texts and images of the retrieval index.
+All runs were performed on a single L40s GPU.
 More details in Appendix B.
-For comparison to medically pre‑trained LVLMs, we include competitive baseline models: BiomedGPT (Luo et al.
-2023); three LLaVA-Med (Li et al.
-2023) variants; two LVLM variants based on PMC CLIP (MedVInT-TE and MedVInT-TD) (Zhang et al.
-2023b); and three variants based on InternVL (Chen et al.
-2024): MedDr, MedDr + RAD (He et al.
-2024a), and GSCo (He et al.
-2024b).
-Full details are in Appendix C.
-Results
-Table 1(a) and Table 1(b) demonstrate the effectiveness of Jomed for two backbone models, Pixtral and Qwen2-vl, across five medical imaging classification benchmarks and three visual question answering benchmarks.
-Jomed achieved superior results compared to both the standard fine-tuned multimodal RAG baseline (trained only with retrieval augmentation fine-tuning, as detailed in Section Reader Retrieval Augmentation Fine-Tuning) and the Reader baseline (LVLM trained without retrieval), for both backbones and all benchmarks.
-In Table 1(a), for classification tasks, Jomed with the Qwen2-vl backbone achieved a mean improvement of 0.05 in F1 and 0.03 in ACC over fine-tuned multimodal RAG, with F1 gains of 0.10 and 0.08 in two datasets.
-Jomed with Pixtral achieved a mean improvement of 0.03 in F1 and ACC.
-In Table 1(b), for VQA tasks, Jomed with Pixtral and Qwen2-vl achieved a mean improvement of 0.02 in ACC (Exact Match) for closed question answering.
-For open question answering, Jomed with Qwen2-vl achieved a mean improvement of 0.03 in token recall, while Jomed with Pixtral achieved an improvement of 0.02.
+Baselines.
+We compare to several RAG baselines.
+RAD He et al.
+(2024a), a retrieval-based approach for classification tasks, where the label of the most similar training image is used as context.
+We integrate RAD with both backbone models (Qwen2-VL and Pixtral).
+MMed-RAG Xia et al.
+(2024a), a recent state-of-the-art multimodal RAG framework, using two backbones: LLaVA-Med and LLaVA, both paired with a medically pre-trained retriever (CLIP).
+We also adopt a standard fine-tuned RAG baseline in which the retriever is fixed and only the reader is fine-tuned, corresponding to the first phase of our approach.
+Finally, we compare our retriever loss with the base Perplexity Distillation Loss (PDist) Izacard et al.
+(2023), integrated into Jomed in place of our customized retrieval loss.
+Full implementation details are in Appendix C.
+For comparison to medically pre‑trained LVLMs, we include competitive baselines: BiomedGPT Luo et al.
+(2023); three LLaVA-Med Li et al.
+(2023) variants; two LVLM variants based on PMC CLIP (MedVInT-TE and MedVInT-TD) Zhang et al.
+(2023b); and three InternVLChen et al.
+(2024) variants: MedDr, MedDr + RAD He et al.
+(2024a), and GSCo He et al.
+(2024b).
+Details in Appendix C.
+4.2 Results
+Table 1(a) and Table 1(b) show the effectiveness of Jomed on five medical classification and three VQA benchmarks.
+We compare with methods of similar computational cost; all LVLMs were not extensively pre-trained on medical data, whereas the MMed-RAG retriever was.
+Jomed consistently outperforms MMed-RAG, RAD, the fine-tuned RAG baseline, Jomed + PDist loss, and the Reader baseline (LVLM without retrieval).
 In addition, we evaluated our model in an oracle scenario.
-Instead of the final fusion of logits (Section Inference), we applied an oracle that, given the model responses for different retrieved images, selects the correct answer if it exists.
+Instead of the final fusion of logits (Section 3.3), we applied an oracle that, given the model responses for different retrieved images, selects the correct answer if it exists.
+[^2] For open-ended questions we do not conduct this analysis as there is no simple boundary between correct/incorrect.
 Our findings in Tables 1(a) and 1(b) show that the correct answer exists in a large proportion of cases, yielding superior results.
@@
 Simply taking the label with highest confidence or label with highest average confidence across retrieved images, underperforms.
-This motivates us to explore the use of the o3 multimodal reasoning model to detect this image (see below).
-Benchmarking versus pretrained LVLMs for classification.
-In Tables 2 and 3, we include previously reported results for several leading medically pretrained LVLM methods (detailed in Section Experimental Setup).
-Jomed with both backbones delivered competitive performance relative to models that underwent extensive pretraining.
-For multi-label classification, Jomed outperformed the pre-trained models.
-For binary and multi-class classification tasks, Jomed with Pixtral matched or exceeded all models except GSCo, which achieved 0.03 higher ACC on the Breast dataset.
-Overall, our models performed strongly across four medical benchmarks compared to medical pretrained models.
-We are not aware of any medical LVLM evaluated on Retina.
-To our knowledge, no other LVLMs have been evaluated on our benchmarks.
-Additionally, in Table 2 we report only ACC, as F1 scores were not reported for the compared models.
-Similarly, in Table 3, we report only F1 scores.
-We also evaluated a medical baseline, Med-Flamingo, with a pretrained medical retriever (BiomedCLIP (Johnson, Douze, and Jégou 2019)).
-We observed that Med-Flamingo generally underperformed, and showed limited improvement from utilizing the retrieved context to enhance performance, which we suspect may be attributed to its relatively older LLaMA backbone (Touvron et al.
-2023).
-We further try BiomedCLIP as a retriever with our general-purpose LVLMs, which showed improvements over FT RAG, with a similar pattern to the general-purpose retriever.
-Full results in Appendix D.
-Benchmarking versus pretrained LVLMs for visual question answering.
-In Table 4, we also compare our approach against medical pretrained LVLM methods (methods detailed in Section Experimental Setup) for VQA.
-Here too, Jomed delivered competitive performance relative to extensively pretrained models.
-For closed question tasks, Jomed with Pixtral and Qwen2-vl was superior to or matched current models.
-For open-ended question answering, Jomed with Qwen2-vl backbone matched LLaVA-MedVicuna on the SLAKE dataset and achieved slightly lower results for PathVQA, showing a 0.01 - 0.02 token recall gap compared to pretrained LVLMs.
-Note that we do not compare our method on VQA-RAD, since our approach is trained on an internal split of training, validation, and test sets, whereas VQA-RAD provides only official training and test splits.
+This motivates us to explore the use of the o3 multimodal reasoning model to detect this image.
+see Section 5.
 Backbone Model Breast Derma Retina VinDr BRSET VQA-RAD SLAKE Mean Mean
@@
 Jomed .91 .87 .85 .67 .62 .53 .71 .10 .64 .42 .75 .91 .75 .52
-Table 5: Analysis of joint training effect: for inconsistent retrieval predictions, we can see that applying joint training boosts performance significantly for all datasets.
+Table 5: Analysis of joint optimization effect: for inconsistent retrieval predictions, we can see that applying joint optimization boosts performance significantly for all datasets.
 For consistent retrieval predictions, the results are slightly improved.
@@
 Pixtral 16% 13% 16% 66% 70% 7% 8%
-Table 6: Proportion of cases with inconsistent predictions, per dataset, for Jomed based on Qwen2-vl and Pixtral.
-Analysis
+Table 6: Proportion inconsistent predictions cases, per dataset, for Jomed based on Qwen2-vl and Pixtral.
+Benchmarking versus medically pretrained LVLMs for classification.
+In Table 3, we include previously reported results for several leading medically pretrained LVLM methods (detailed in Section 4.1).
+Jomed demonstrates competitive performance compared to models that underwent extensive medical pretraining (pretraining sizes summarized in Table 2).
+Jomed with the Pixtral backbone matches or exceeds all models except GSCo on the Breast dataset.
+To our knowledge, no medically pretrained LVLMs have been evaluated on the Retina benchmark, and we are not aware of any prior LVLM results on our other classification benchmarks.
+For binary and multi-class classification, we report only accuracy (ACC), as prior works did not report F1 scores.
+Conversely, for multi-label classification, we report only F1 scores.
+We also evaluated a medical baseline, Med-Flamingo, which generally underperformed and showed limited benefit from retrieval augmentation.
+We attribute this to its relatively older LLaMA backbone Touvron et al.
+(2023).
+Additionally, we experimented with using BiomedCLIP Johnson et al.
+(2019b) as the retriever for our general-purpose LVLMs, which exhibited a similar trend to the non-medical retriever.
+Full results in Appendix D.3.
+Benchmarking against pretrained LVLMs for visual question answering.
+In Table 4, we compare our approach with medically pretrained LVLM methods (detailed in Section 4.1) for VQA.
+Jomed achieves competitive performance relative to extensively pretrained models.
+For closed-question tasks, Jomed outperforms or matches existing models.
+For open-ended question answering, Jomed with the Qwen2-vl backbone matches LLaVA-MedVicuna on the SLAKE dataset and LLaVA-MedLLaVA on the PathVQA dataset.
+Note that we do not report results on VQA-RAD, since our method was trained using an internal split of training and validation, whereas VQA-RAD provides only official training and test splits.
+5 Analysis
 Jomed boosts empirically challenging cases.
@@
 Overall, Jomed improves performance on inconsistent retrieval predictions by +0.12 in accuracy and +0.13 in F1 score when using the Qwen2‑vl backbone, and by +0.09 in both accuracy and F1 score with the Pixtral backbone, while also offering a slight improvement on the consistent.
-We also observed that inconsistent retrieval predictions are empirically more challenging for all models (both with and without retrieval).
-There is approximately a 10–20 point performance gap between the inconsistent and consistent.
+We also observed that inconsistent retrieval predictions are empirically more challenging for the reader, FT RAG and Jomed models, with a 10–20 point performance gap between the inconsistent and consistent.
 Performance with o3 as reranker.
-We have observed that results for inconsistent predictions remain relatively low (Table 5), even after applying joint training (Table 5).
+Results for inconsistent predictions remain relatively low (Table 5), even after applying joint optimization (Table 5).
 However, in the oracle analysis reported above, we found that Jomed can sometimes retrieve images that lead to correct predictions, together with other images that lead to wrong predictions.
-We thus explore whether a state-of-the-art multimodal model, when used as an optional reranker, could close the gap toward oracle performance for inconsistent retrieval predictions.
-Specifically, we investigated whether the o3 reasoning model (OpenAI 2025) could effectively select the image from the retrieved set that offers the most predictive information, guiding the model toward correct classification.
+We thus explore whether a state-of-the-art multimodal model, when used as an optional reranker, could close the gap toward oracle performance.
+Specifically, we investigated whether the o3 reasoning model OpenAI (2025) could effectively select the image from the retrieved set that offers the most predictive information, guiding the model toward correct classification.
 We provided o3 with an image to classify, along with four retrieved images+captions, and tasked it with identifying the one containing the most predictive information.
-(detailed prompt in Appendix D)
-Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in Jomed.
-Thus, we were unable to bridge the gap to oracle performance, and conclude that reranking in this setting remains a significant challenge even for a frontier model.
-We report full results in Appendix D.
+Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in Jomed.We conclude that reranking in this setting remains a significant challenge even for a frontier model.
+Results in Appendix D.2.
 Additional analyses and ablations.
-We include several robustness ablations and analyses in Appendix D.
-We observed that our model is robust to scenarios where the retrieval is noisy or not given to the models.
-Further, we observed that the retrieval itself includes predictive knowledge without being given the query.
-details in Appendix D.
-Conclusion
-We explored the effectiveness of a multimodal retriever jointly optimized with a general-purpose LVLM.
-Our lightweight approach yielded substantial improvements, including in challenging cases involving inconsistent retrieval predictions.
+We evaluated the contribution of each component in our method in Appendix D.1.
+Additionally, we present several robustness ablations in Appendix D.4.
+We demonstrate robustness across different scenarios.
+6 Conclusion
+We demonstrated that joint optimization of general-purpose LVLMs and multimodal retrievers can achieve competitive medical diagnosis performance through lightweight fine-tuning, without medical pre-training.
+Future work may explore whether this approach generalizes as a cost-effective alternative to domain-specific pre-training when task-specific data and domain knowledge bases are available but large-scale pre-training is prohibitive.
+In addition, we identified and substantially mitigated inconsistent retrieval predictions, where different retrieved candidates lead to conflicting diagnoses.
 Our oracle analysis revealed a considerable performance gap between actual results and what was theoretically achievable using the retrieved images, providing a foundation for future research on closing this gap.
+Future work may also look into the prevalence and impact of inconsistent retrieval predictions more broadly.
+Limitations
+Sequential rather than end-to-end optimization.
+We acknowledge that, in our method, gradients do not flow simultaneously between the retriever and reader modules.
+Nevertheless, the reader leverages signals from the retriever during training to improve its performance, while the retriever, in turn, benefits from feedback provided by the reader to refine its retrieved content.
+This bidirectional interaction effectively induces a joint optimization effect within our framework.
+Scope of evaluation.
+Our evaluation focuses on classification and visual question-answering tasks.
+We did not evaluate our method on report generation, which requires different capabilities (longer-form generation, clinical writing conventions) and would benefit from dedicated investigation.
+Additionally, while our method shows consistent improvements across diverse imaging modalities (ultrasound, fundus photography, X-ray, histopathology, dermatoscopy, CT) and tasks (visual question answering, malignant lesion classification, diabetic retinopathy severity grading), evaluation on additional specialized modalities (e.g., PET, Microscopy, OCT) and tasks (e.g., organ classification, retinal OCT disease classification) would further validate generalizability.
 Appendix A Dataset
+A.1 Dataset for Evaluation
+Our datasets include real-world hospital datasets (BRSET Nakayama et al.
+(2024) and VinDr‑PCXR Pham et al.
+(2022)) alongside a variety of classification and visual question answering datasets.
+We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA).
+Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive.
+Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed Casey (2022).
+Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study.
+Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge to compensate for sparse training samples.
+For the classification tasks, we used the following datasets:
+BreastMNIST (nickname: Breast) Al-Dhabyani et al.
+(2020) is a binary classification dataset of breast ultrasound.
+Following the official split, we use 546 for training, 78 for validation and 156 for testing.
+RetinaMNIST (nickname: Retina) Liu et al.
+(2022) is a multi-label classification dataset of retina Fundus Camera.
+Following the official split, we use 1,080 for training, 120 for validation and 400 for testing.
+DermaMNIST (nickname: Derma) Tschandl et al.
+(2018) is a multi-class classification dataset of common pigmented skin lesions.
+Following the official split, we use 7,007 for training, 1,003 for validation and 2,005 for testing.
+The dataset contains clinical images from 7 different diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis, dermatofibroma, melanoma, melanocytic nevi, and vascular lesions.
+VinDr-PCXR Pham et al.
+(2022) is a large, open pediatric chest X-ray dataset collected in Vietnam (2020–2021) containing 9,125 posteroanterior scans from patients under 10 years old.
+It provides both lesion-level bounding-box annotations for 36 findings and image-level labels for multi-label of 15 diagnoses, curated by experienced radiologists.
+We follow the official split, which includes 7,728 training and 1,397 test images, with an additional internal split of the training data into 85% for training and 15% for validation.
+BRSET Nakayama et al.
+(2024) is a Brazilian multilabel ophthalmological dataset comprising retinal fundus images annotated with 14 distinct pathological findings.
+The dataset, composed of 16,266 images, presents a challenging real-world scenario for multilabel classification.
+There is no publicly available split; therefore, we apply an internal split for train, validation, and test sets (70% train, 10% validation, and 20% test).
+Breast, Derma, and Retina taken from the large-scale MNIST-like collection of standardized biomedical images, including 12 datasets for 2D and 6 datasets for 3D.
+We especially used MedMNIST+ Yang et al.
+(2023), which is a higher resolution extension of the original MedMNIST.
+We used the highest resolution of 224 × 224.
+For medical question answering, we have three well-known datasets:
+VQA-RAD Lau et al.
+(2018) is a clinician-annotated visual question answering dataset focused on radiology images (e.g., X-ray, CT, MRI).
+It pairs each image with both open-ended and yes/no questions.
+Following the official split, we use 1,753 questions for training and 453 for testing.
+In addition, we further partition the training set into 85% for training and 15% for validation.
+PathVQA He et al.
+(2020) is a pathology-focused medical VQA dataset built from textbooks and the PEIR digital library, comprising 4,289 images and 32,632 question–answer pairs spanning both open-ended and yes/no questions.
+We follow the official split provided by the authors: 19,700 for training, 6,260 for validation, and 6,720 for testing.
+SLAKE Liu et al.
+(2021) is a bilingual medical VQA dataset.
+We use the English subset, SLAKE-English, which comprises 642 radiology images and 7.03k English QA pairs.
+We follow the official train/validation/test split of 4.92k/1.05k/1.06k QA pairs, and include all question types, covering both open-ended and closed-ended formats.
+A.2 Dataset for Index
+For construction of the Index we used three large datasets of (image, text) pairs: PMC-OA Lin et al. (2023), ROCO Rückert et al. (2024) and MIMIC-CXR Johnson et al. (2019a):
+PMC-OA: is a large-scale dataset that contains 1.65M image-text pairs. The figures and captions from PubMed Central, 2,478,267 available papers are covered.
+ROCO: is an image-caption dataset collected from PubMed. It filters out all the compound or non-radiological images, and consists of 81K samples.
+MIMIC-CXR: is the largest chest X-ray dataset, containing 377,110 samples (image-report pairs). Each image is paired with a clinical report describing findings from doctors.
 Backbone Retrieval type Breast Derma Retina VQA-RAD Mean
@@
 We evaluate on Breast, Derma, Retina, and the closed-question subset of VQA-RAD.
-Dataset for Evaluation
-Our datasets include real-world hospital datasets (BRSET (Nakayama et al.
-2024) and VinDr‑PCXR (Pham, Tran, and Nguyen 2022)) alongside a variety of classification and visual question answering datasets.
-We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA).
-Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive.
-Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed (Casey 2022).
-Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study.
-Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge to compensate for sparse training samples.
-For the classification tasks, we used the following datasets:
-BreastMNIST (nickname: Breast) (Al-Dhabyani et al.
-2020) is a binary classification dataset of breast ultrasound.
-Following the official split, we use 546 for training, 78 for validation and 156 for testing.
-RetinaMNIST (nickname: Retina) (Liu et al.
-2022) is a multi-label classification dataset of retina Fundus Camera.
-Following the official split, we use 1,080 for training, 120 for validation and 400 for testing.
-DermaMNIST (nickname: Derma) (Tschandl, Rosendahl, and Kittler 2018) is a multi-class classification dataset of common pigmented skin lesions.
-Following the official split, we use 7,007 for training, 1,003 for validation and 2,005 for testing.
-The dataset contains clinical images from 7 different diagnostic categories: actinic keratoses, basal cell carcinoma, benign keratosis, dermatofibroma, melanoma, melanocytic nevi, and vascular lesions.
-VinDr-PCXR (Pham, Tran, and Nguyen 2022) is a large, open pediatric chest X-ray dataset collected in Vietnam (2020–2021) containing 9,125 posteroanterior scans from patients under 10 years old.
-It provides both lesion-level bounding-box annotations for 36 findings and image-level labels for multi-label of 15 diagnoses, curated by experienced radiologists.
-We follow the official split, which includes 7,728 training and 1,397 test images, with an additional internal split of the training data into 85% for training and 15% for validation.
-BRSET (Nakayama et al.
-2024) is a Brazilian multilabel ophthalmological dataset comprising retinal fundus images annotated with 14 distinct pathological findings.
-The dataset, composed of 16,266 images, presents a challenging real-world scenario for multilabel classification.
-There is no publicly available split; therefore, we apply an internal split for train, validation, and test sets (70% train, 10% validation, and 20% test).
-Breast, Derma, and Retina taken from the large-scale MNIST-like collection of standardized biomedical images, including 12 datasets for 2D and 6 datasets for 3D.
-We especially used MedMNIST+ (Yang et al.
-2023), which is a higher resolution extension of the original MedMNIST.
-We used the highest resolution of 224 × 224.
-For medical question answering, we have three well-known datasets:
-VQA-RAD (Lau et al.
-2018) is a clinician-annotated visual question answering dataset focused on radiology images (e.g., X-ray, CT, MRI).
-It pairs each image with both open-ended and yes/no questions.
-Following the official split, we use 1,753 questions for training and 453 for testing.
-In addition, we further partition the training set into 85% for training and 15% for validation.
-PathVQA (He et al.
-2020) is a pathology-focused medical VQA dataset built from textbooks and the PEIR digital library, comprising 4,289 images and 32,632 question–answer pairs spanning both open-ended and yes/no questions.
-We follow the official split provided by the authors: 19,700 for training, 6,260 for validation, and 6,720 for testing.
-SLAKE (Liu et al.
-2021) is a bilingual medical VQA dataset.
-We use the English subset, SLAKE-English, which comprises 642 radiology images and 7.03k English QA pairs.
-We follow the official train/validation/test split of 4.92k/1.05k/1.06k QA pairs, and include all question types, covering both open-ended and closed-ended formats.
-Dataset for Index
-For construction of the Index we used three large datasets of (image, text) pairs: PMC-OA (Lin et al. 2023), ROCO (Rückert et al. 2024) and MIMIC-CXR (Johnson et al. 2019):
-PMC-OA: is a large-scale dataset that contains 1.65M image-text pairs. The figures and captions from PubMed Central, 2,478,267 available papers are covered.
-ROCO: is an image-caption dataset collected from PubMed. It filters out all the compound or non-radiological images, and consists of 81K samples.
-MIMIC-CXR: is the largest chest X-ray dataset, containing 377,110 samples (image-report pairs). Each image is paired with a clinical report describing findings from doctors.
 Backbone Model Breast Derma Retina
@@
 Appendix B Implementation Details
-Retrieval Augmentation
+B.1 Retrieval Augmentation
 As described earlier, for each image–question pair, we retrieve $r$ image–text pairs, with $r=4$.
-We use the input image as the query and Jina-CLIP (Xiao, Mastrapas, and Wang 2024) as the retriever head in our multimodal retriever.
-The index is constructed with FAISS (Johnson, Douze, and Jégou 2019) over MIMIC-CXR, PMC-OA, and ROCO, which are fully described in Section A in Appendix A.
+We use the input image as the query and Jina-CLIP Xiao et al.
+(2024) as the retriever head in our multimodal retriever.
+The index is constructed with FAISS Johnson et al.
+(2019b) over MIMIC-CXR, PMC-OA, and ROCO, which are fully described in Section A.2 in Appendix A.
 To embed images, we use the Jina-CLIP visual head (the same model used for retrieval); to embed captions/reports, we use the Jina-CLIP text head.
 The index is stored in float16 due to storage constraints.
-We also experimented with BiomedCLIP (Zhang et al.
-2023a).
-Reader Fine-Tuning
-We use Pixtral (12B) (Agrawal et al.
-2024) and Qwen2-vl (7B) (Wang et al.
-2024) as base models.
+We also experimented with BiomedCLIP Zhang et al.
+(2023a).
+B.2 Reader Fine-Tuning
+We use Pixtral (12B) Agrawal et al.
+(2024) and Qwen2-vl (7B) Wang et al.
+(2024) as base models.
 We train the models with and without retrieval augmentation to assess their effect.
 The retrieval-augmented prompt is described in Appendix E.
-Note, we also tried Med-Flamingo (9B) (Moor et al.
-2023) as a base model.
-We fine-tune Pixtral and Qwen2-vl using LLaMA-Factory (Zheng et al.
-2024) on a single NVIDIA L40 (48 GB) GPU for 10 epochs.
-We use a learning rate of $2\times 10^{-5}$ and apply Low-Rank Adaptation (LoRA) (Hu et al.
-2021) for parameter-efficient fine-tuning.
+Note, we also tried Med-Flamingo (9B) Moor et al.
+(2023) as a base model.
+We fine-tune Pixtral and Qwen2-vl using LLaMA-Factory Zheng et al.
+(2024) on a single NVIDIA L40 (48 GB) GPU for 10 epochs.
+We use a learning rate of $2\times 10^{-5}$ and apply Low-Rank Adaptation (LoRA) Hu et al.
+(2021) for parameter-efficient fine-tuning.
 To all models (including baseline models), we conduct a grid search over batch size (2, 4, 6) and whether to freeze the vision head, selecting the best configuration by validation performance.
@@
 We use a learning rate of $2\times 10^{-5}$.
-Multimodal Retriever Fine-Tuning
+B.3 Multimodal Retriever Fine-Tuning
 We train the retrieval model to fetch relevant context while keeping the reader frozen.
@@
 In our experiments, retrieving more candidates further improved retriever performance.
-Evaluation Metrics
+B.4 Evaluation Metrics
 Classification.
@@
 For open-ended questions, we compute token-level Recall and F1 between the predicted token set $\hat{A}$ and the reference token set $A$:
-$\mathrm{Prec}=\frac{|\hat{A}\cap A|}{|\hat{A}|},\quad\mathrm{Rec}=\frac{|\hat{A}\cap A|}{|A|},\quad\mathrm{F1}=\frac{2\,\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}.$
+$\mathrm{Prec}=\frac{|\hat{A}\cap A|}{|\hat{A}|},\quad\mathrm{Rec}=\frac{|\hat{A}\cap A|}{|A|},\quad$
+$\mathrm{F1}=\frac{2\,\mathrm{Prec}\cdot\mathrm{Rec}}{\mathrm{Prec}+\mathrm{Rec}}.$
 For closed-ended questions, we use Exact Match.
-Backbone
-BiomedCLIP
-Variant
-Breast Derma Retina Vindr BREST VQA-RAD SLAKE PathVQA Mean
+Model
+Breast Derma Retina Vindr BREST VQARAD SLAKE PathVQA Mean
 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC ACC ACC ACC F1
-Qwen2-vl FT RAG .84 .81 .73 .38 .63 .45 .54 .09 .49 .27 .77 .89 .92 .72 .40
-Jomed .87 .82 .75 .40 .65 .50 .58 .15 .50 .39 .79 .91 .92 .74 .45
-Pixtral FT RAG .88 .84 .79 .57 .57 .46 .55 .09 .44 .31 .74 .87 .88 .72 .45
-Jomed .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
+FT RAGQwen2-vl .84 .81 .73 .38 .63 .45 .54 .09 .49 .27 .77 .89 .92 .72 .40
+JomedQwen2-vl .87 .82 .75 .40 .65 .50 .58 .15 .50 .39 .79 .91 .92 .74 .45
+FT RAGPixtral .88 .84 .79 .57 .57 .46 .55 .09 .44 .31 .74 .87 .88 .72 .45
+JomedPixtral .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
 Table 10: Performance comparison of Jomed against FT RAG baseline across medical imaging classification and VQA tasks.
@@
 VQA results are reported for the closed-question subset using exact match accuracy.
-MedFlamingo
-Variant
+Model
 Breast Derma Retina VQARAD
 ACC F1 ACC F1 ACC F1 ACC
-Reader only .82 .77 .67 .54 .57 .40 .65
+Reader
+only
+.82 .77 .67 .54 .57 .40 .65
 Jomed .81 .77 .68 .54 .60 .51 .69
@@
 Appendix C Baseline Methods
+C.1 RAG baseline
+We evaluate our method against several representative RAG-based and multimodal retrieval approaches to ensure a fair and comprehensive comparison.
+RAD He et al.
+(2024a) is a retrieval-based method designed for classification tasks.
+During inference, the most similar image from the training set is retrieved, and its supervised label is directly used as the prediction.
+We integrate RAD with both of our backbone models (Qwen2-VL and Pixtral) under the same data splits and retrieval settings.
+MMed-RAGXia et al.
+(2024a) represents a recent state-of-the-art multimodal RAG approach in which the retriever and LVLM are trained independently.
+We follow the official GitHub implementation, first training the CLIP module and then performing DPO training using two backbone models: LLaVA-Med and LLaVA.
+Fusion-in-Decoder (FiD) Pipeline Izacard and Grave (2020) serves as a standard RAG-style baseline.
+In this setting, the retriever is frozen, and only the reader is fine-tuned while processing retrieved image–caption pairs as contextual input.
+Perplexity Distillation Loss (PDist) Izacard et al.
+(2023).
+For each query and candidate document, we compute the reduction in perplexity (or equivalently, the increase in likelihood) of the correct output when the language model is conditioned on that document.
+From these likelihoods, we derive a “posterior” over documents (proportional to their contributions), and train the retriever to mimic that posterior via a KL divergence objective.
+In doing so, the retriever is guided to prioritize documents that most effectively support the language model’s generation, thereby aligning retrieval with generation quality.
+C.2 Medical Pre - trained baseline models
 We benchmark against state-of-the-art medical large vision–language models (LVLMs) that underwent large-scale medical pre-training:
 BiomedGPT.
-An open multimodal generative pre-trained transformer for biomedicine that aligns diverse biological/biomedical modalities with natural language; we include it as a strong domain baseline (Luo et al.
-2023).
+An open multimodal generative pre-trained transformer for biomedicine that aligns diverse biological/biomedical modalities with natural language; we include it as a strong domain baseline Luo et al.
+(2023).
 LLaVA-Med.
-A biomedical adaptation of LLaVA trained via curriculum (biomedical figure–caption alignment followed by instruction tuning); we evaluate three commonly used releases as separate baselines (Li et al.
-2023).
+A biomedical adaptation of LLaVA trained via curriculum (biomedical figure–caption alignment followed by instruction tuning); we evaluate three commonly used releases as separate baselines Li et al.
+(2023).
 MedVInT-TE and MedVInT-TD.
-Two implementations of Medical Visual Instruction Tuning from PMC-VQA (Zhang et al.
-2023b).
-Both adopt a PMC-CLIP vision encoder (Lin et al.
-2023).
+Two implementations of Medical Visual Instruction Tuning from PMC-VQA Zhang et al.
+(2023b).
+Both adopt a PMC-CLIP vision encoder Lin et al.
+(2023).
 TE uses an encoder-style text pathway feeding a multimodal decoder, while TD concatenates text tokens with visual features to a decoder-only pathway; we report both as distinct baselines.
 InternVL-based baselines: MedDr, MedDr+RAD, and GSCo.
-MedDr is a generalist medical VLLM trained on large-scale instruction-style data curated from diagnosis-guided bootstrapping and medical image descriptions, covering multiple modalities and tasks (He et al.
-2024a).
-MedDr+RAD augments MedDr at inference with Retrieval-Augmented Diagnosis (RAD): for a test image, similar cases are retrieved and summarized as contextual guidance in the prompt (He et al.
-2024b).
-GSCo (Generalist–Specialist Collaboration) is a two-stage framework that (i) builds a generalist GFM (MedDr) and lightweight specialist models, and (ii) performs collaborative inference via Mixture-of-Expert Diagnosis (MoED; using specialist predictions as references) and RAD (using specialists to retrieve similar cases) (He et al.
-2024b).
-GSCo evaluates across a large multi-dataset benchmark; we include GSCo as a strong InternVL-based baseline along with its MedDr and MedDr+RAD components (He et al.
-2024b; Chen et al.
-2024).
+MedDr is a generalist medical VLLM trained on large-scale instruction-style data curated from diagnosis-guided bootstrapping and medical image descriptions, covering multiple modalities and tasks He et al.
+(2024a).
+MedDr+RAD augments MedDr at inference with Retrieval-Augmented Diagnosis (RAD): for a test image, similar cases are retrieved and summarized as contextual guidance in the prompt He et al.
+(2024b).
+GSCo (Generalist–Specialist Collaboration) is a two-stage framework that (i) builds a generalist GFM (MedDr) and lightweight specialist models, and (ii) performs collaborative inference via Mixture-of-Expert Diagnosis (MoED; using specialist predictions as references) and RAD (using specialists to retrieve similar cases) He et al.
+(2024b).
+GSCo evaluates across a large multi-dataset benchmark; we include GSCo as a strong InternVL-based baseline along with its MedDr and MedDr+RAD components He et al.
+(2024b); Chen et al.
+(2024).
 Appendix D Additional Analysis
-Performance with a State-of-the-Art LVLM Reranker
+D.1 Analysis of model components
+We analyze the contribution of each part in the model in Table 7(a) and in Table 7(b).
+The result of the first stage of training - the reader retrieval augmentation fine-tuning is the FT RAG then we check the contribution of training only the text retriever head and then finally is the contribution of training both the text and image retriever head as Jomed.
+We also explore the utility of converting the open question into a closed one in our retrieval loss.
+If we don’t apply o3 model to convert the open question, we only trained on the closed one, which is referred to in Table 7(b) as Closed Question Only.
+D.2 Performance with a State-of-the-Art LVLM Reranker
 We evaluate whether a state-of-the-art large vision–language model (LVLM), used as an optional reranker, can improve performance on inconsistent-retrieval predictions.
-Specifically, we test the o3 reasoning model (OpenAI 2025) as a reranker: given a query image to classify and four retrieved image–caption pairs, o3 is asked to select the single pair that is most informative for predicting the correct label.
+Specifically, we test the o3 reasoning model OpenAI (2025) as a reranker: given a query image to classify and four retrieved image–caption pairs, o3 is asked to select the single pair that is most informative for predicting the correct label.
 We compare this reranking to four alternatives: (i) using the top-1 retrieved candidate (no reranking), (ii) selecting the model’s prediction with the highest confidence (maximum logit), (iii) using o3 with captions only (no images), and (iv) the Jomed aggregation strategy.
@@
 See results in Table 9.
-Jomed variants with medical pre-trained backbones
+D.3 Jomed variants with medical pre-trained backbones
 Although our work focuses on general-purpose LVLMs, we also evaluated a medical-specific baseline.
@@
 The performance gap between Med-Flamingo and our reader approach was small—0.01 in accuracy and 0.03 in F1.
-We hypothesize that this narrow gap may be due to Med-Flamingo’s relatively older LLaMA backbone (Touvron et al.
-2023).
+We hypothesize that this narrow gap may be due to Med-Flamingo’s relatively older LLaMA backbone Touvron et al.
+(2023).
 Results are presented in Table 11.
@@
 Full results are presented in Table 10.
-Retrieval Robustness Analyses.
+D.4 Retrieval Robustness Analyses.
 We strive for our model to be robust in cases where the retrieval is noisy.
@@
 Appendix E Prompt Design
-Here, we detail the prompts used in our experiments and analysis.
-To explore the utility of o3 as re-rankers of the most predictive (image, caption) pair, as described in Section D in Appendix D, we provide both models with a detailed prompt for the task, shown in Prompt 1.
-Additionally, we present the prompts used to train Jomed for each of the datasets: Prompts 2, 4, 3, 6, 5, and the prompt for visual question answering tasks Prompt 7.
-1 Task:
-2
-3 You are an expert assistant selecting the most informative reference for medical image classification.
-4
-5 Description:
-6 A patients [MRI/CT/X-ray] image needs to be classified into one of the following categories: [List of possible labels].
-7 You are given four candidate reference pairs, each consisting of a caption and its associated image, drawn from PubMed literature.
-8 Your goal is to determine which single candidate provides the most useful information to help an AI model correctly classify the provided image.
-9
-10 Input Information:
-11 - Medical image: The patients MRI/CT/X-ray image is provided here.
-12 - Classification task: Classify the image into one of the following categories: [List of possible labels].
-13 - Candidates:
-14 - Candidate 0:
-15 - Caption: [Caption 0]
-16 - Image: [Image 0]
-17 - Candidate 1:
-18 - Caption: [Caption 1]
-19 - Image: [Image 1]
-20 - Candidate 2:
-21 - Caption: [Caption 2]
-22 - Image: [Image 2]
-23 - Candidate 3:
-24 - Caption: [Caption 3]
-25 - Image: [Image 3]
-26
-27 Instructions:
-28 1. For each candidate, carefully assess how informative its caption and image are for the current classification task.
-29 2. Specifically, evaluate whether the candidate describes imaging features, findings, or clinical context relevant to distinguishing among the listed categories.
-30 3. Compare all candidates and select the one that gives the clearest, most discriminative information to aid the classification.
-31 4. Output only the number of the selected candidate in the following format:
-32 Caption Number: [your selected number]
-Listing 1: Prompt used for o3
+We detail the prompts used in our experiments and analyses.
+To convert open questions into closed ones with the o3 model, we use the prompt in Prompt 1.
+To assess the utility of o3 as a re-ranker of the most predictive image–caption pair (Section D.2, Appendix D), we provide both LVLM backbones (Pixtral and Qwen2-VL) with the task-specific prompt in Prompt 2.
+We also include the training prompts for Jomed across datasets—Prompts 3, 5, 4, 7, 6—and the VQA prompt (Prompt 8).
+Convert the following open-ended question into a closed yes/no question based on the given answer.
+The new question should be answerable with "{expected_answer}".
+Original Question: {question}
+Original Answer: {answer}
+Expected New Answer: {expected_answer}
+Please provide only the new closed question without any additional text or explanation.
+Listing 1: Prompt used for o3 to convert open question into closed one
+Task:
+You are an expert assistant selecting the most informative reference for medical image classification.
+Description:
+A patients [MRI/CT/X-ray] image needs to be classified into one of the following categories: [List of possible labels].
+You are given four candidate reference pairs, each consisting of a caption and its associated image, drawn from PubMed literature.
+Your goal is to determine which single candidate provides the most useful information to help an AI model correctly classify the provided image.
+Input Information:
+- Medical image: The patients MRI/CT/X-ray image is provided here.
+- Classification task: Classify the image into one of the following categories: [List of possible labels].
+- Candidates:
+- Candidate 0:
+- Caption: [Caption 0]
+- Image: [Image 0]
+- Candidate 1:
+- Caption: [Caption 1]
+- Image: [Image 1]
+- Candidate 2:
+- Caption: [Caption 2]
+- Image: [Image 2]
+- Candidate 3:
+- Caption: [Caption 3]
+- Image: [Image 3]
+Instructions:
+1. For each candidate, carefully assess how informative its caption and image are for the current classification task.
+2. Specifically, evaluate whether the candidate describes imaging features, findings, or clinical context relevant to distinguishing among the listed categories.
+3. Compare all candidates and select the one that gives the clearest, most discriminative information to aid the classification.
+4. Output only the number of the selected candidate in the following format:
+Caption Number: [your selected number]
+Listing 2: Prompt used for o3
 […]
-Listing 2: Prompt used by Jomed for Breast classification
+Listing 3: Prompt used by Jomed for Breast classification
 […]
-Listing 3: Prompt used by Jomed for Retina classification
+Listing 4: Prompt used by Jomed for Retina classification
 […]
@@
 […]
-Listing 4: Prompt used by Jomed for Derma classification
+Listing 5: Prompt used by Jomed for Derma classification
 […]
@@
 […]
-Listing 5: Prompt used by Jomed for Vindr-PCXR classification
+Listing 6: Prompt used by Jomed for Vindr-PCXR classification
 […]
@@
 […]
-Listing 6: Prompt used by Jomed for BRSET classification
+Listing 7: Prompt used by Jomed for BRSET classification
 […]
-Listing 7: Prompt used by Jomed for visual question answering tasks
+Listing 8: Prompt used by Jomed for visual question answering tasks

### v4

@@
-Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis
+LVLM-Aware Multimodal Retrieval for RAG-Based Medical Diagnosis with General-Purpose Models
 Abstract
-Retrieving relevant visual and textual information from medical literature and hospital records can enhance diagnostic accuracy for clinical image interpretation.
-We develop a multimodal retrieval model jointly optimized with an LVLM for medical diagnosis, unlike standard RAG which doesn’t backpropagate LVLM errors to the retriever.
-Using only general-purpose backbones with lightweight fine-tuning, our model achieves competitive results with medically-pretrained models on clinical classification and VQA tasks.
-In a novel analysis, we find that different top-retrieved images often yield different predictions for the same target, and that these cases are challenging for all models, even for non-retrieval models.
-Our joint retrieval optimization significantly improves these cases over standard RAG.
-However, oracle analysis reveals that while the correct diagnosis is frequently achievable using one of the top retrieved images, in practice there is a large performance gap from the oracle, and rerankers using frontier LVLMs do not close this gap—leaving ample room for improvement by future methods.
-Code available at https://github.com/Nirmaz/JOMED.
-Lightweight Joint Optimization of General-Purpose Vision-Language Models and Retrievers for RAG-Based Medical Diagnosis
+Retrieving visual and textual information from medical literature and hospital records can enhance diagnostic accuracy for clinical image interpretation.
+However, multimodal retrieval-augmented diagnosis is highly challenging.
+We explore a lightweight mechanism for enhancing diagnostic performance of retrieval-augmented LVLMs.
+We train a lightweight LVLM-aware multimodal retriever, such that the retriever learns to return images and texts that guide the LVLM toward correct predictions.
+In our low-resource setting, we perform only lightweight fine-tuning with small amounts of data, and use only general-purpose backbone models, achieving competitive results in clinical classification and VQA tasks compared to medically pre-trained models with extensive training.
+In a novel analysis, we highlight a previously unexplored class of errors that we term inconsistent retrieval predictions: cases where different top-retrieved images yield different predictions for the same target.
+We find that these cases are challenging for all models, even for non-retrieval models, and that our retrieval optimization mechanism significantly improves these cases over standard RAG.
+However, our analysis also sheds light on gaps in the ability of LVLMs to utilize retrieved information for clinical predictions.
+Code and models available at: https://github.com/Nirmaz/JOMED.
+LVLM-Aware Multimodal Retrieval for RAG-Based Medical Diagnosis with General-Purpose Models
 Nir Mazor1, Tom Hope1,2 1School of Computer Science and Engineering, The Hebrew University of Jerusalem 2The Allen Institute for AI (AI2)
 1 Introduction
-Figure 1: We jointly optimize a multimodal retriever and a Large Vision-Language Model for medical tasks.
+Figure 1: We optimize a multimodal retriever and a Large Vision-Language Model for medical tasks.
 We achieve competitive results without resource-intensive medical pre-training and significantly improve performance on challenging cases where different retrieved images lead to inconsistent retrieval prediction.
@@
 (2025).
-In this work, we explore a lightweight fine-tuning approach for taking a general-purpose LVLM and optimizing it together with a general-purpose multimodal retriever for medical diagnosis tasks.
-Different from the common multimodal RAG setup, we perform joint optimization of the LVLM and retriever.
-This teaches the LVLM to utilize retrieved knowledge—images, captions and clinical reports from both literature and hospital records—while at the same time tuning the retriever to find information that leads to correct predictions by the LVLM.
+In this work, we explore a lightweight fine-tuning approach using a general-purpose LVLM with a general-purpose multimodal retriever for medical diagnosis tasks.
+Unlike standard multimodal RAG, we train an LVLM-aware multimodal retriever to find knowledge—clinical images, captions, and reports from both medical literature and hospital records—that leads to correct LVLM predictions.
 Poorly optimized retrieval mechanisms can mislead models Yoran et al.
 (2023); Sun et al.
-(2024); as shown in Figure 2, a RAG model incorrectly diagnoses a benign ultrasound when using a cancerous retrieved image, but joint optimization leads to correct retrieval and prediction.
+(2024).
+As shown in Figure 2, a RAG model incorrectly classifies a benign ultrasound image when conditioned on a cancerous retrieved image, whereas our LVLM-aware retrieval optimization leads to the correct predicted diagnosis.
 Our method involves sequential multimodal training with a dual-head retriever architecture and a customized retrieval loss and visual question answering (VQA) training recipe.
@@
 Importantly, our focus in this work is not to chase SOTA results but to shed light on several interesting observations.
-One, that a simple and effective joint LVLM-retriever optimization mechanism can lead to a substantial boost in results over current multimodal RAG methods and should thus be more widely adopted for medical diagnosis.
+One, that a simple and effective LVLM-aware retrieval optimization mechanism can lead to a substantial boost in results over current multimodal RAG methods and should thus be more widely adopted for medical diagnosis; beyond downstream performance, our analysis also shows improved relevance of retrieval candidates.
 Second, that this mechanism achieves competitive results by using models with no medical pre-training (for neither the LVLM nor the retriever) and only lightweight fine-tuning.
@@
 (2024) and has potentially important implications considering the resource-intensive nature of pre-training processes.
-Third, previous work on joint optimization in the general domain required large-scale pre-training followed by few-shot task-specific adaptation Izacard et al.
+Third, related work in the general domain that tuned both retrieval and generation models required large-scale pre-training followed by few-shot task-specific adaptation Izacard et al.
 (2023); Hu et al.
@@
 (2024).
-In contrast, we conduct joint optimization directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples).
-This provides first evidence about the feasibility and utility of lightweight data-efficient joint optimization directly on downstream tasks as opposed to in the pre-training setting.
-We are also the first to explore and show the utility of joint optimization in the multimodal medical domain.
+In contrast, we conduct our optimization directly on downstream tasks with no pretraining and only lightweight fine-tuning (as few as 546 samples).
+This provides first evidence about the feasibility and utility of lightweight data-efficient generation-aware retrieval optimization directly on downstream tasks as opposed to in the pre-training setting.
+In addition to their reliance on extensive training, another gap in this line of work is that the various techniques developed in the LLM community for tuning retrievers to align with LLMs Izacard et al.
+(2023); Shi et al.
+(2023); Lin et al.
+(2024) have yet to be explored in the multimodal setting with LVLMs.
 Fourth and importantly, we conduct a novel analysis that finds that our method helps in particular to address a class of errors we term inconsistent retrieval predictions.
@@
 (2022).
-As we show, our joint optimization mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases.
+As we show, our retrieval mechanism significantly mitigates this issue and achieves large improvement over standard fine-tuned RAG on these cases.
 We further discover that for a large proportion of these inconsistent cases, at least one retrieved candidate enables the model to predict the correct answer in an oracle setting in which we provide the model with the correct retrieved candidate.
 This suggests useful information is being retrieved but is often lost among less helpful candidates.
-To address this, we explored using the powerful multimodal model o3 OpenAI (2025) as a reranker to identify the most predictive instances.
+To address this, we explored using the powerful o3 model OpenAI (2025) as a reranker to identify predictive candidates.
 We found it did not close the performance gap to the oracle and was comparable or inferior to our method, leaving significant room for future improvement.
 Our contributions are three-fold:
-  • We jointly optimize a multimodal retriever and an LVLM for medical classification and VQA.
-We use only general-purpose backbones without any medical pretraining, and show that by retrieving medical information and co-optimizing the retriever and LVLM we can close much of the gap to expensive medically pre-trained models.
-  • Unlike previous general-domain joint optimization that required pre-training, we conduct joint optimization directly on downstream tasks with only lightweight data-efficient fine-tuning.
-  • We identify inconsistent retrieval predictions, challenging cases where different retrieved images lead to different predictions for the same query.
-Our joint optimization method significantly improves performance on these cases when compared to standard RAG.
-Figure 2: joint optimization impacts inconsistent retrieval predictions.
-After joint optimization, retrieved images show greater alignment with query image labels.
-In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after joint optimization, the retrieved image is less directly related to the wrong label (c).
+  • We perform LVLM-aware retrieval optimization for a multimodal retriever and an LVLM on medical classification and VQA tasks (including text generation VQA).
+We use only general-purpose backbones without medical pretraining, and show that by retrieving relevant medical evidence and training the retriever to guide the LVLM, we close much of the gap to expensive medically pre-trained models and achieve superior performance compared to general-purpose RAG methods.
+  • Unlike previous general-domain methods that required pre-training, we conduct lightweight data-efficient optimization of a retriever and LVLM directly on downstream tasks.
+  • We identify inconsistent retrieval predictions, challenging cases where retrieved images lead to different predictions for the same query.
+Our method significantly improves performance on these cases over standard RAG.
+Figure 2: LVLM-aware multimodal retrieval finetuning impacts inconsistent retrieval predictions.
+After LVLM-aware multimodal retrieval finetuning, retrieved images show greater alignment with query image labels.
+In Breast, a cancer-free ultrasound query (a) initially retrieved a lesion image (b), however, after finetuning, the retrieved image is less directly related to the wrong label (c).
 In VQA-RAD, retrieval shifted from an unrelated medical condition (e) to an image depicting the condition of the query image.
 2 Related Work
-Retrieval Augmented joint optimization has been explored primarily in unimodal (text only) NLP with work such as ATLAS Izacard et al.
+Retrieval-Augmented LLM-Aware Optimization has been explored primarily in unimodal (text only) NLP with work such as ATLAS Izacard et al.
 (2023), which conducted large-scale pretraining of both reader and retriever models and evaluated in zero-shot and few-shot scenarios in the general domain.
 REVEAL Hu et al.
-(2023) extended this to the multimodal setting, achieving state-of-the-art results in visual tasks in the general domain with extensive pre-training.
+(2023) extended this to the multimodal setting in the general domain with extensive pre-training; REVEAL used an encoder-decoder architecture with a T5 generator—to our knowledge, since REVEAL no work has explored optimizing multimodal retrievers and generators with causal decoder models and modern VLMs.
 Shi et al.
-(2023) proposed training textual retrievers based on reader performance, influencing subsequent joint-optimization methods.
+(2023) proposed training textual retrievers based on reader performance, influencing subsequent retrieval alignment LLM-aware optimization methods.
 Lin et al.
-(2024) introduced another joint optimization variant and proposed sequential training: first training the reader with retrieval augmentation, then training the retriever.
+(2024) introduced another LLM-aware optimization variant and proposed first training the reader with retrieval augmentation, then training the retriever.
 Siriwardhana et al.
-(2023) demonstrated joint optimization’s effectiveness for domain adaptation.
-Our work differs by being the first to explore joint optimization optimized directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training.
-Our extensive novel analyses are the first to shed light on inconsistent predictions and how joint optimization on downstream tasks helps address them.
+(2023) demonstrated LLM-aware optimization’s effectiveness for domain adaptation.
+Our work differs by being the first to explore generator-aware retrieval optimization directly for downstream tasks, with minimal fine-tuning, as opposed to requiring large-scale pre-training.
+Prior work, such as ATLAS and REVEAL, follows a two-step approach—first conducting large-scale pre-training, then evaluating in few-shot or zero-shot settings, while we use a lightweight one-step approach instead.
+Our method includes a new loss and new training methodologies.
+Our novel analyses are the first to shed light on inconsistent predictions and methods to help address them.
 Multimodal Retrieval Augmentation in Medical Applications.
@@
 (2023), which integrated retrieved text and images with query modalities.
-With Large Vision-Language Models (LVLMs), He et al.
-(2024a) proposed retrieving relevant labeled examples during inference.
+With LVLMs, He et al.
+(2024a) proposed retrieving labeled examples during inference.
 Xia et al.
-(2024b) addressed LVLMs’ misalignment with medical knowledge using retrieval-augmented generation to fetch similar medical reports, employing contrastive learning and context selection strategies.
+(2024b) improved LVLM RAG with contrastive learning and context selection strategies.
 Xia et al.
-(2024a) proposed domain-aware retrieval for diverse medical data sources, while Wu et al.
+(2024a) proposed domain-aware retrieval for diverse medical data, while Wu et al.
 (2025) combined retrieval augmentation with knowledge graphs.
+All current multimodal medical RAG methods are not LVLM-aware, and the retriever is trained disjointly from the LVLM.
 3 Method
-Figure 3: The two-phase Jomed training.
+Figure 3: The two-phase CLARE training.
 The LVLM is first trained with a frozen retriever on augmented prompts containing retrieved image–caption/report pairs.
@@
 Overview.
-In this section, we present our methodology, termed Jomed (Joint Optimization of Multimodal Retrieval for Enhanced Medical Diagnosis).
+In this section, we present our methodology, termed CLARE (Clinical LVLM-Aware Retrieval).
 Our task involves medical image classification and visual question answering.
-Jomed comprises two main components: a multimodal retriever and a reader.
+CLARE comprises two main components: a multimodal retriever and a reader.
 Given an input patient image and a diagnostic question, the multimodal retriever identifies relevant medical knowledge—images and their associated captions or hospital reports—that provide predictive information.
 The reader then analyzes the retrieved candidates along with the question to generate an answer.
-We jointly optimize the reader to better leverage the retrieved information and the multimodal retriever to enhance its ability to select informative candidates for the reader.
+We optimize the reader to better leverage the retrieved information and the multimodal retriever to enhance its ability to select informative candidates for the reader.
 Figure 3 illustrates our framework.
@@
 (2024a); Xia et al.
-(2024b, a) our method does not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip Xiao et al.
+(2024b, a) we do not use medical pretraining, and instead we use readily available general LVLMs (Pixtral and Qwen2-vl) and a general retriever, jina-clip Xiao et al.
 (2024).
-The selected LVLMs can process multiple images and texts, allowing us to incorporate both the retrieved image and the retrieved caption/report into the model context.
-We also compare our results to Med-Flamingo Moor et al.
-(2023).
-Med-Flamingo is a medical LVLM which is also able to process multiple images, unlike other medical LVLM available at the time of conducting our experiments.
+The selected LVLMs can process multiple images and texts, allowing us to incorporate both the retrieved image and the retrieved text.
+We also compare to Med-Flamingo Moor et al.
+(2023), a medical LVLM which is also able to process multiple images, unlike other medical LVLMs available at the time of conducting our experiments.
 We now elaborate on the details.
-3.1 Reader Retrieval Augmentation Fine-Tuning
+3.1 Reader Fine-Tuning
 We first fine-tune the reader with retrieval augmentation.
@@
 where $\theta$ represents the parameters of the LVLM, and $\circ$ denotes the concatenation operation.
-3.2 Sequential Multimodal Retriever Fine-Tuning
-Our dual-head multimodal retriever has a text-based head for retrieving relevant (image, text) pairs based on textual similarity to the query, and an image-based head for retrieving pairs based on visual similarity.
-To train this retriever, we adopt a sequential training strategy, first optimizing the text retrieval head followed by the image retrieval head.
-During this process, the reader model remains frozen, focusing solely on improving the retriever’s embedding space.
+3.2 Sequential Retriever Fine-Tuning
+Our dual-head multimodal retriever has a text-based head for retrieving relevant (image, text) pairs based on textual similarity to the query, and an image-based head for visual similarity.
+We adopt a sequential training strategy, first optimizing the text retrieval head followed by the image head.
+During this process, the reader remains frozen, solely on improving retriever’s embedding space.
 Similarly to Izacard et al.
@@
 Let $z_{k}=(\mathbf{i}_{k},\mathbf{t}_{k})$ denote the $k$-th retrieved image-text pair and $q=(\mathbf{i},\mathbf{q})$ denote the query pair.
-The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved pair $z_{k}$ and query $q$:
-$p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q),$
-where $p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q)$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
+The posterior distribution reflects the model’s confidence in predicting the correct answer $\mathbf{a}$ given a retrieved pair $z_{k}$ and query $q$: $p_{k}\propto p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q)$, where $p_{\text{LVLM}}(\mathbf{a}\mid z_{k}\circ q)$ is the probability assigned by the LVLM to $\mathbf{a}$, and $\circ$ denotes prepending the retrieved chunk to the query.
 We sharpen the LVLM’s output distribution by restricting it to only the relevant class tokens from the benchmark’s possible answers.
@@
 $p_{\text{LVLM}}(a\mid q)=\sum_{k=1}^{K}p_{\text{LVLM}}(a\mid z_{k}\circ q)\cdot p_{\text{R}}(z_{k}\mid q),$
-where $p_{\text{R}}(z_{k}\mid q)$ is:
-$p_{\text{R}}(z_{k}\mid q)=\frac{\exp(s(q,z_{k}))}{\sum_{j=1}^{K}\exp(s(q,z_{j}))},$
-and $s(q,z_{k})$ is the similarity score between the query and the retrieved chunk.
+where $p_{\text{R}}(z_{k}\mid q)=\frac{\exp(s(q,z_{k}))}{\sum_{j=1}^{K}\exp(s(q,z_{j}))}$, and $s(q,z_{k})$ is the similarity score between the query and the retrieved chunk.
 Backbone (Size) Model Breast Derma Retina VinDr-PCXR BRSET Mean
@@
 FT RAG .85 .82 .71 .42 .62 .48 .55 .09 .48 .27 .64 .42
-Jomed + PDist .86 .83 .73 .47 .63 .49 .57 .09 .47 .29 .65 .43
-Jomed .87 .84 .76 .50 .65 .50 .57 .14 .49 .37 .67 .47
-Jomedoracle .87 .84 .85 .64 .69 .51 .64 .14 .52 .41 .71 .51
+CLARE + PDist .86 .83 .73 .47 .63 .49 .57 .09 .47 .29 .65 .43
+CLARE .87 .84 .76 .50 .65 .50 .57 .14 .49 .37 .67 .47
+CLAREoracle .87 .84 .85 .64 .69 .51 .64 .14 .52 .41 .71 .51
 Pixtral (12B) Reader .82 .77 .75 .52 .55 .44 .48 .08 .44 .33 .61 .43
@@
 FT RAG .88 .85 .79 .60 .57 .47 .49 .09 .47 .33 .64 .47
-Jomed + PDist .88 .84 .80 .62 .60 .47 .50 .09 .45 .35 .65 .47
-Jomed .90 .87 .80 .62 .60 .51 .56 .14 .51 .37 .67 .50
-Jomedoracle .93 .90 .83 .67 .63 .54 .67 .15 .57 .41 .73 .53
+CLARE + PDist .88 .84 .80 .62 .60 .47 .50 .09 .45 .35 .65 .47
+CLARE .90 .87 .80 .62 .60 .51 .56 .14 .51 .37 .67 .50
+CLAREoracle .93 .90 .83 .67 .63 .54 .67 .15 .57 .41 .73 .53
 Backbone (Size) Model VQA-RAD SLAKE PathVQA Mean
@@
 FT RAG .76 .45 .88 .81 .91 .33 .85 .53
-Jomed + PDist .77 .45 .89 .81 .92 .32 .86 .53
-Jomed .79 .48 .90 .84 .93 .38 .87 .57
-Jomedoracle .86 - .92 - .95 - .91 -
+CLARE + PDist .77 .45 .89 .81 .92 .32 .86 .53
+CLARE .79 .48 .90 .84 .93 .38 .87 .57
+CLAREoracle .86 - .92 - .95 - .91 -
 Pixtral (12B) Reader .72 .38 .83 .80 .87 .26 .81 .48
 FT RAG .74 .41 .88 .81 .88 .31 .83 .51
-Jomed + PDist .75 .41 .89 .81 .89 .32 .84 .51
-Jomed .76 .45 .90 .84 .90 .36 .85 .55
-Jomedoracle .88 - .92 - .95 - .92 -
-Table 1: Results for classification (a) and VQA (b) with non-medically pretrained LVLM backbones (see Tables 3, 4 for comparisons to with medical pre-training; MMed-RAG’s retrieval component is medically pre-trained).
-joint optimization consistently leads to the best results.
+CLARE + PDist .75 .41 .89 .81 .89 .32 .84 .51
+CLARE .76 .45 .90 .84 .90 .36 .85 .55
+CLAREoracle .88 - .92 - .95 - .92 -
+Table 1: Results for classification (a) and VQA (b) with non-medically pretrained LVLM backbones (see Tables 3, 4 for comparisons to medical pre-training; MMed-RAG’s retrieval component is medically pre-trained).
+CLARE consistently leads to the best results.
 For VQA, we evaluate closed questions using the exact match metric and open questions using token recall.
-Using an oracle reranker (for classification and closed VQA; §4.2) demonstrates that even larger improvements are achievable with Jomed’s top-retrieved images in some datasets.
-MMed-RAG
-LLaVaMed
-variants
-MedDr
-variants
-GSCo
-MedVInT
-variants
-Medical Pre-Training
-Size
-600K 600K 255K 255K 177K
-Table 2: Medical pre-training size for LVLMs under extensive medical pretraining.
+Using an oracle reranker (for classification and closed VQA; §4.2) demonstrates that even larger improvements are achievable with CLARE’s top-retrieved images in some datasets.
+MMed-RAG LLaVaMed variants MedDr variants GSCo MedVInT variants
+Medical Pre-Training Size 600K 600K 255K 255K 177K
+Table 2: Medical pre-training sizes of existing LVLMs.
 4 Experiments
@@
 GSCoInternVL .93 - .09 .33
-Jomed Qwen2-vl .87 .76 .14 .37
-JomedPixtral .90 .80 .14 .37
+CLARE Qwen2-vl .87 .76 .14 .37
+CLAREPixtral .90 .80 .14 .37
 Table 3: Comparison to prior reported results of medical pre-trained models, for binary and multiclass classification (Breast and Derma; accuracy), and for multi-label classification (VinDr-PCXR and BRSET; F1).
@@
 LLaVA-MedBioMedCLIP .87 .87 .91 .39
-JomedQwen2-vl .90 .84 (.84) .93 .38 (.37)
-JomedPixtral .90 .84 (.82) .90 .36 (.35)
+CLAREQwen2-vl .90 .84 (.84) .93 .38 (.37)
+CLAREPixtral .90 .84 (.82) .90 .36 (.35)
 Table 4: Comparison to medical pre-trained models for visual question answering. We report the token recall metric, reported for LLaVA-Med variants and token F1 reported for BiomedGPT (in red).
@@
 (2024) and VinDr‑PCXR Pham et al.
-(2022)) alongside a variety of classification and visual question answering datasets.
+(2022)) alongside a variety of classification and VQA datasets.
 We focus on a low-resource data-efficient setting (training sets ranging from 546 to 7007 samples in classification, 1790-19,700 in VQA).
 Medical image annotation is a resource-intensive task, demanding expert annotators whose availability is limited and expensive.
-Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for the specific context in which they are deployed Casey (2022).
+Clinical AI also often has poor generalization across heterogeneous medical centers and patient populations, often requiring that models be trained for each deployment Casey (2022).
 Such constraints often restrict researchers to datasets comprising only a few thousand samples for a given study.
 Retrieval augmentation is well-suited for this setting, as it is known to benefit low‑data regimes the most by leveraging external knowledge.
-We consider three types of classification tasks: binary, multi-class, and multi-label.
+We consider binary, multi-class, and multi-label classification.
 These include BreastMNIST (“Breast”; breast ultrasound imaging, binary) Al-Dhabyani et al.
 (2020), DermaMNIST (“Derma”; pigmented skin lesion images, multi-class) Tschandl et al.
-(2018), RetinaMNIST (“Retina”; retinal fundus images with multi-class disease annotations) Liu et al.
-(2022), and two challenging real-world multi-label datasets: VinDr-PCXR (a chest X-ray dataset with 15 labels) Pham et al.
-(2022) and BRSET (an ophthalmology dataset with 14 labels) Nakayama et al.
+(2018), RetinaMNIST (“Retina”; retinal fundus images, multi-class) Liu et al.
+(2022), and two challenging real-world multi-label datasets: VinDr-PCXR (chest X-rays, 15 labels) Pham et al.
+(2022) and BRSET (ophthalmology, 14 labels) Nakayama et al.
 (2024).
-For visual question answering, we use widely adopted benchmarks: VQA‑RAD Lau et al.
+For VQA, we use widely adopted benchmarks: VQA‑RAD Lau et al.
 (2018), SLAKE‑English Liu et al.
@@
 RAD He et al.
-(2024a), a retrieval-based approach for classification tasks, where the label of the most similar training image is used as context.
-We integrate RAD with both backbone models (Qwen2-VL and Pixtral).
+(2024a), a retrieval-based approach for classification, where the label of the most similar training image is used as context; we use RAD with both Qwen2-VL and Pixtral.
 MMed-RAG Xia et al.
-(2024a), a recent state-of-the-art multimodal RAG framework, using two backbones: LLaVA-Med and LLaVA, both paired with a medically pre-trained retriever (CLIP).
+(2024a), a recent state-of-the-art multimodal RAG framework in which the retriever is optimized independently from the LVLM and is not aware of the LVLM during training.
+We evaluated MMed-RAG using backbones LLaVA-Med and LLaVA, both paired with a medically pre-trained retriever (CLIP).
 We also adopt a standard fine-tuned RAG baseline in which the retriever is fixed and only the reader is fine-tuned, corresponding to the first phase of our approach.
 Finally, we compare our retriever loss with the base Perplexity Distillation Loss (PDist) Izacard et al.
-(2023), integrated into Jomed in place of our customized retrieval loss.
-Full implementation details are in Appendix C.
-For comparison to medically pre‑trained LVLMs, we include competitive baselines: BiomedGPT Luo et al.
+(2023), integrated into CLARE in place of our customized loss.
+Implementation details are in Appendix C.
+To compare to medically pre‑trained LVLMs, we include competitive baselines: BiomedGPT Luo et al.
 (2023); three LLaVA-Med Li et al.
@@
 (2024) variants: MedDr, MedDr + RAD He et al.
-(2024a), and GSCo He et al.
+(2024a), GSCo He et al.
 (2024b).
@@
 4.2 Results
-Table 1(a) and Table 1(b) show the effectiveness of Jomed on five medical classification and three VQA benchmarks.
+Table 1(a) and Table 1(b) show the effectiveness of CLARE on five medical classification and three VQA benchmarks.
 We compare with methods of similar computational cost; all LVLMs were not extensively pre-trained on medical data, whereas the MMed-RAG retriever was.
-Jomed consistently outperforms MMed-RAG, RAD, the fine-tuned RAG baseline, Jomed + PDist loss, and the Reader baseline (LVLM without retrieval).
+CLARE consistently outperforms MMed-RAG, RAD, the fine-tuned RAG baseline, CLARE + PDist loss, and the Reader baseline (LVLM without retrieval).
 In addition, we evaluated our model in an oracle scenario.
@@
 Our findings in Tables 1(a) and 1(b) show that the correct answer exists in a large proportion of cases, yielding superior results.
-This shows that in these cases, Jomed is able to surface in its top retrieved images an image which could lead to a correct prediction, however the information gets lost in the process of prediction fusion with other retrieved images.
-Simply taking the label with highest confidence or label with highest average confidence across retrieved images, underperforms.
-This motivates us to explore the use of the o3 multimodal reasoning model to detect this image.
-see Section 5.
+This shows that in these cases, CLARE is able to surface in its top retrieved images an image which could lead to a correct prediction, however the information gets lost in the process of prediction fusion with other retrieved images.
+Simply taking the label with highest confidence or label with highest average confidence, underperforms.
+This motivates us to explore the o3 multimodal reasoning model to detect this image (Section 5).
 Backbone Model Breast Derma Retina VinDr BRSET VQA-RAD SLAKE Mean Mean
-ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
-Closed
-Acc
-Closed
-Acc
-ACC F1
+ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 Closed Acc Closed Acc ACC F1
 Qwen2-vl Inconsistent Reader .40 .40 .50 .25 .35 .31 .41 .07 .33 .21 .43 .70 .40 .25
 RAG .40 .40 .52 .36 .40 .28 .46 .07 .49 .25 .33 .84 .45 .27
-Jomed .80 .80 .62 .46 .52 .29 .48 .16 .50 .36 .47 .86 .58 .41
+CLARE .80 .80 .62 .46 .52 .29 .48 .16 .50 .36 .47 .86 .58 .41
 Consistent Reader .84 .77 .83 .29 .63 .45 .56 .09 .34 .22 .77 .85 .64 .36
 RAG .86 .82 .90 .49 .66 .50 .66 .10 .43 .23 .82 .89 .70 .43
-Jomed .88 .84 .91 .50 .66 .55 .66 .10 .41 .37 .82 .91 .70 .47
+CLARE .88 .84 .91 .50 .66 .55 .66 .10 .41 .37 .82 .91 .70 .47
 Pixtral Inconsistent Reader .60 .59 .45 .33 .40 .32 .38 .07 .46 .28 .66 .64 .46 .32
 RAG .72 .71 .44 .45 .35 .37 .38 .08 .39 .26 .72 .65 .46 .37
-Jomed .84 .84 .50 .50 .42 .46 .49 .13 .45 .35 .77 .68 .54 .46
+CLARE .84 .84 .50 .50 .42 .46 .49 .13 .45 .35 .77 .68 .54 .46
 Consistent Reader .86 .80 .80 .56 .56 .44 .66 .08 .61 .37 .74 .85 .70 .45
 RAG .91 .87 .84 .67 .62 .50 .71 .08 .64 .38 .75 .91 .74 .50
-Jomed .91 .87 .85 .67 .62 .53 .71 .10 .64 .42 .75 .91 .75 .52
-Table 5: Analysis of joint optimization effect: for inconsistent retrieval predictions, we can see that applying joint optimization boosts performance significantly for all datasets.
+CLARE .91 .87 .85 .67 .62 .53 .71 .10 .64 .42 .75 .91 .75 .52
+Table 5: Analysis of LVLM-aware multimodal retrieval fine-tuning effect: for inconsistent retrieval predictions, we can see that applying LVLM-aware multimodal retrieval fine-tuning boosts performance significantly for all datasets.
 For consistent retrieval predictions, the results are slightly improved.
@@
 Pixtral 16% 13% 16% 66% 70% 7% 8%
-Table 6: Proportion inconsistent predictions cases, per dataset, for Jomed based on Qwen2-vl and Pixtral.
+Table 6: Proportion inconsistent predictions cases, per dataset, for CLARE based on Qwen2-vl and Pixtral.
 Benchmarking versus medically pretrained LVLMs for classification.
-In Table 3, we include previously reported results for several leading medically pretrained LVLM methods (detailed in Section 4.1).
-Jomed demonstrates competitive performance compared to models that underwent extensive medical pretraining (pretraining sizes summarized in Table 2).
-Jomed with the Pixtral backbone matches or exceeds all models except GSCo on the Breast dataset.
+In Table 3, we include previously reported results for leading medically pretrained LVLM methods (detailed in Section 4.1).
+CLARE demonstrates results competitive with models that underwent extensive medical pretraining (pretraining sizes in Table 2).
+CLARE with the Pixtral backbone matches or exceeds all models except GSCo on the Breast dataset.
 To our knowledge, no medically pretrained LVLMs have been evaluated on the Retina benchmark, and we are not aware of any prior LVLM results on our other classification benchmarks.
@@
 Additionally, we experimented with using BiomedCLIP Johnson et al.
-(2019b) as the retriever for our general-purpose LVLMs, which exhibited a similar trend to the non-medical retriever.
+(2019b) as the retriever for our general-purpose LVLMs, which exhibited a similar trend.
 Full results in Appendix D.4.
@@
 In Table 4, we compare our approach with medically pretrained LVLM methods (detailed in Section 4.1) for VQA.
-Jomed achieves competitive performance relative to extensively pretrained models.
-For closed-question tasks, Jomed outperforms or matches existing models.
-For open-ended question answering, Jomed with the Qwen2-vl backbone matches LLaVA-MedVicuna on the SLAKE dataset and LLaVA-MedLLaVA on the PathVQA dataset.
+CLARE achieves competitive performance relative to extensively pretrained models.
+For closed-question tasks, CLARE outperforms or matches existing models.
+For open-ended question answering, CLARE with the Qwen2-vl backbone matches LLaVA-MedVicuna on the SLAKE dataset and LLaVA-MedLLaVA on the PathVQA dataset.
 Note that we do not report results on VQA-RAD, since our method was trained using an internal split of training and validation, whereas VQA-RAD provides only official training and test splits.
 5 Analysis
-Jomed boosts empirically challenging cases.
-We observed that the performance gap is substantially larger for predictions classified as inconsistent retrieval predictions (the proportion of inconsistent retrieval predictions is reported in Table 6).
-We define an inconsistent retrieval prediction as one in which the model predicts different labels for retrieved candidates given the same target query (see Figure 1).
-Specifically, if the model’s predictions vary across retrieved candidates, the instance is categorized as an inconsistent retrieval prediction.
-Conversely, if the model consistently predicts the same label for all retrieved candidates, it is classified as a consistent retrieval prediction.
-Both the consistent and inconsistent retrieval prediction sets are established after the initial reader training stage, and their performance is assessed following joint multimodal retriever fine-tuning.
+CLARE boosts empirically challenging cases.
+We observe that the performance gap is substantially larger for predictions classified as inconsistent retrieval predictions (Table 6 shows the proportion of such cases).
+We define an inconsistent retrieval prediction as one in which the model predicts different labels for retrieved candidates given the same query (see Figure 1), i.e., if the model’s predictions vary across retrieved candidates.
+Conversely, if the model predicts the same label for all retrieved candidates, the instance is consistent.
+Consistent and inconsistent prediction sets are established after the initial reader training stage, and their performance is assessed following LVLM-aware multimodal retrieval fine-tuning.
 Detailed results are presented in Table 5.
-Overall, Jomed improves performance on inconsistent retrieval predictions by +0.12 in accuracy and +0.13 in F1 score when using the Qwen2‑vl backbone, and by +0.09 in both accuracy and F1 score with the Pixtral backbone, while also offering a slight improvement on the consistent.
-We also observed that inconsistent retrieval predictions are empirically more challenging for the reader, FT RAG and Jomed models, with a 10–20 point performance gap between the inconsistent and consistent.
-Performance with o3 as reranker.
-Results for inconsistent predictions remain relatively low (Table 5), even after applying joint optimization (Table 5).
-However, in the oracle analysis reported above, we found that Jomed can sometimes retrieve images that lead to correct predictions, together with other images that lead to wrong predictions.
+Overall, CLARE improves performance on inconsistent retrieval predictions by +0.12 in accuracy and +0.13 in F1 score when using the Qwen2‑vl backbone, and by +0.09 in both accuracy and F1 score with the Pixtral backbone, while also offering a slight improvement on the consistent.
+We also observed that inconsistent retrieval predictions are empirically more challenging for the reader, FT RAG and CLARE models, with a 10–20 point performance gap.
+Performance with reranking.
+Results for inconsistent predictions remain relatively low (Table 5).
+However, in the oracle analysis reported above, we found that CLARE can sometimes retrieve images that lead to correct predictions, together with other images that lead to wrong predictions.
 We thus explore whether a state-of-the-art multimodal model, when used as an optional reranker, could close the gap toward oracle performance.
-Specifically, we investigated whether the o3 reasoning model OpenAI (2025) could effectively select the image from the retrieved set that offers the most predictive information, guiding the model toward correct classification.
+Specifically, we investigated whether the o3 reasoning model OpenAI (2025) could select a retrieved image that leads to correct prediction.
 We provided o3 with an image to classify, along with four retrieved images+captions, and tasked it with identifying the one containing the most predictive information.
-Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in Jomed.We conclude that reranking in this setting remains a significant challenge even for a frontier model.
+Our results show that o3’s performance generally surpasses simply taking the image with highest confidence, but o3 is generally inferior compared to fusing logits in CLARE.
+We conclude that reranking in this setting remains a significant challenge even for a frontier model.
 Results in Appendix D.3.
-Additional analyses and ablations.
-We evaluated the contribution of each component in our method in Appendix D.1.
-Additionally, we present several robustness ablations in Appendix D.4.
-We demonstrate robustness across different scenarios.
+Retrieval relevance and additional analyses.
+We explore the utility of our method in retrieving more relevant candidates.
+In our setting where no retrieval ground-truth labels are available, we use GPT-5.2 as a judge to compare relevance of retrieved candidates before and after the LVLM-aware retriever fine-tuning.
+Our method improves candidate relevance with an average 17% win rate, as detailed in Appendix D.1.
+We further evaluate the contribution of each component in Appendix D.2 and present robustness ablations in Appendix D.5.
 6 Conclusion
-We demonstrated that joint optimization of general-purpose LVLMs and multimodal retrievers can achieve competitive medical diagnosis performance through lightweight fine-tuning, without medical pre-training.
+We demonstrated that LVLM-aware multimodal retrieval achieves competitive medical diagnosis performance through lightweight, data-efficient fine-tuning, without medical pre-training.
 Future work may explore whether this approach generalizes as a cost-effective alternative to domain-specific pre-training when task-specific data and domain knowledge bases are available but large-scale pre-training is prohibitive.
@@
 Limitations
-Sequential rather than end-to-end optimization.
-We acknowledge that, in our method, gradients do not flow simultaneously between the retriever and reader modules.
-Nevertheless, the reader leverages signals from the retriever during training to improve its performance, while the retriever, in turn, benefits from feedback provided by the reader to refine its retrieved content.
-This bidirectional interaction effectively induces a joint optimization effect within our framework.
 Scope of evaluation.
-Our evaluation focuses on classification and visual question-answering tasks.
+Our evaluation focuses on classification and visual question-answering tasks, including both open questions (a text-generation setting) and closed questions (closed-form answer selection).
 We did not evaluate our method on report generation, which requires different capabilities (longer-form generation, clinical writing conventions) and would benefit from dedicated investigation.
@@
 o3 multi-image .77 .80 .43 .33 .43 .35
-Jomed .77 .80 .57 .43 .51 .47
+CLARE .77 .80 .57 .43 .51 .47
 Pixtral Top-1 .72 .75 .46 .48 .29 .33
@@
 o3 multi-image .73 .73 .43 .42 .34 .37
-Jomed .76 .77 .45 .47 .34 .37
+CLARE .76 .77 .45 .47 .34 .37
 Table 9: Study of o3 as a reranker.
@@
 We use a learning rate of $2\times 10^{-5}$.
-B.3 Multimodal Retriever Fine-Tuning
+B.3 LVLM-Aware Multimodal Retrieval Fine-Tuning
 We train the retrieval model to fetch relevant context while keeping the reader frozen.
@@
 For closed-ended questions, we use Exact Match.
-Model
-Breast Derma Retina Vindr BREST VQARAD SLAKE PathVQA Mean
+Model Breast Derma Retina Vindr BREST VQARAD SLAKE PathVQA Mean
 ACC F1 ACC F1 ACC F1 ACC F1 ACC F1 ACC ACC ACC ACC F1
 FT RAGQwen2-vl .84 .81 .73 .38 .63 .45 .54 .09 .49 .27 .77 .89 .92 .72 .40
-JomedQwen2-vl .87 .82 .75 .40 .65 .50 .58 .15 .50 .39 .79 .91 .92 .74 .45
+CLAREQwen2-vl .87 .82 .75 .40 .65 .50 .58 .15 .50 .39 .79 .91 .92 .74 .45
 FT RAGPixtral .88 .84 .79 .57 .57 .46 .55 .09 .44 .31 .74 .87 .88 .72 .45
-JomedPixtral .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
-Table 10: Performance comparison of Jomed against FT RAG baseline across medical imaging classification and VQA tasks.
-Both Qwen2-VL and Pixtral backbones show consistent improvements with Jomed, particularly in F1 scores.
+CLAREPixtral .91 .87 .80 .59 .59 .49 .56 .15 .50 .39 .77 .89 .89 .74 .50
+Table 10: Performance comparison of CLARE against FT RAG baseline across medical imaging classification and VQA tasks.
+Both Qwen2-VL and Pixtral backbones show consistent improvements with CLARE, particularly in F1 scores.
 Results demonstrate that BiomedCLIP-based retrieval enhances performance similarly to the general-purpose retriever.
 VQA results are reported for the closed-question subset using exact match accuracy.
-Model
-Breast Derma Retina VQARAD
+Model Breast Derma Retina VQARAD
 ACC F1 ACC F1 ACC F1 ACC
-Reader
-only
-.82 .77 .67 .54 .57 .40 .65
-Jomed .81 .77 .68 .54 .60 .51 .69
-Table 11: Med-Flamingo comparison with and without Jomed retrieval augmentation across medical imaging tasks.
-Jomed achieves best performance on 4 out of 7 metrics, with notable improvements over Reader Only in Retina classification (ACC: 0.57→0.60, F1: 0.40→0.51) and VQA-RAD (0.65→0.69).
+Reader only .82 .77 .67 .54 .57 .40 .65
+CLARE .81 .77 .68 .54 .60 .51 .69
+Table 11: Med-Flamingo comparison with and without CLARE retrieval augmentation across medical imaging tasks.
+CLARE achieves best performance on 4 out of 7 metrics, with notable improvements over Reader Only in Retina classification (ACC: 0.57→0.60, F1: 0.40→0.51) and VQA-RAD (0.65→0.69).
 However, the overall performance gains are more modest than those observed with other backbone architectures, indicating that Med-Flamingo may have limited capacity to leverage additional retrieval context.
@@
 (2024).
+Backbone Winner Breast Derma Retina VinDr BRSET VQA-RAD SLAKE PathVQA Mean
+Qwen2-vl Tie .20 .59 .23 .25 .50 .24 .50 .20 .34
+FT RAG Wins .30 .09 .32 .36 .24 .25 .16 .35 .26
+CLARE Wins .50 .30 .44 .46 .25 .50 .34 .45 .41
+Pixtral Tie .24 .51 .51 .47 .35 .12 .15 .29 .33
+FT RAG Wins .26 .12 .15 .19 .40 .32 .37 .30 .26
+CLARE Wins .50 .35 .34 .34 .52 .54 .47 .41 .43
+Table 12: Analysis of relevance of retrieval candidates before and after LVLM-aware multimodal retrieval fine-tuning.
+Training the multimodal retrieval significantly improves performance, with a mean improvement of 0.17 across Qwen2-vl and Pixtral.
 Appendix D Additional Analysis
-D.1 Analysis of model components
+D.1 Retrieval Relevancy Analysis
+We explore whether LVLM-aware multimodal retrieval fine-tuning improves the quality of retrieval candidates compared to those produced before fine-tuning.
+Importantly, our setup does not include ground-truth retrieval labels; instead, the retriever learns which candidates better steer the frozen LVLM toward the correct downstream prediction.
+To measure candidate improvement, we compare the retriever’s image and text heads before vs.
+after fine-tuning in a head-to-head ranking task.
+We then prompt GPT-5.2 to judge which candidate more closely resembles the ground-truth prediction, assuming that higher similarity to the ground-truth prediction indicates better evidence steering.
+When both candidates are identical, we record a tie.
+To avoid any location biases, we conduct shuffling between entering the model and reducing cost; we use a subset of 10% of the samples.
+Prompt 2 details the prompt for the task.
+Table 12 shows that LVLM-aware multimodal retrieval fine-tuning improves candidate relevance, with a mean gain of 0.17 for both Qwen2-VL and Pixtral, demonstrating the practical utility of our method.
+D.2 Analysis of Model Components
 We analyze the contribution of each part in the model in Table 7(a) and in Table 7(b).
-The result of the first stage of training - the reader retrieval augmentation fine-tuning is the FT RAG then we check the contribution of training only the text retriever head and then finally is the contribution of training both the text and image retriever head as Jomed.
+The result of the first stage of training - the reader retrieval augmentation fine-tuning is the FT RAG then we check the contribution of training only the text retriever head and then finally is the contribution of training both the text and image retriever head as CLARE.
 We also explore the utility of converting the open question into a closed one in our retrieval loss.
 If we don’t apply o3 model to convert the open question, we only trained on the closed one, which is referred to in Table 7(b) as Closed Question Only.
-D.2 Performance with a State-of-the-Art LVLM Reranker
+D.3 Performance with a State-of-the-Art LVLM Reranker
 We evaluate whether a state-of-the-art large vision–language model (LVLM), used as an optional reranker, can improve performance on inconsistent-retrieval predictions.
 Specifically, we test the o3 reasoning model OpenAI (2025) as a reranker: given a query image to classify and four retrieved image–caption pairs, o3 is asked to select the single pair that is most informative for predicting the correct label.
-We compare this reranking to four alternatives: (i) using the top-1 retrieved candidate (no reranking), (ii) selecting the model’s prediction with the highest confidence (maximum logit), (iii) using o3 with captions only (no images), and (iv) the Jomed aggregation strategy.
-Our results show that o3 generally outperforms the maximum-logit baseline but remains inferior to Jomed’s logit-fusion aggregation.
+We compare this reranking to four alternatives: (i) using the top-1 retrieved candidate (no reranking), (ii) selecting the model’s prediction with the highest confidence (maximum logit), (iii) using o3 with captions only (no images), and (iv) the CLARE aggregation strategy.
+Our results show that o3 generally outperforms the maximum-logit baseline but remains inferior to CLARE’s logit-fusion aggregation.
 We evaluated performance on the Breast, Derma, and Retina datasets.
@@
 See results in Table 9.
-D.3 Jomed variants with medical pre-trained backbones
+Backbone Model Breast Derma Retina VinDr BRSET
+ACC F1 ACC F1 ACC F1 ACC F1 ACC F1
+Qwen2-vl CLARE N=2 .87 .83 .74 .50 .63 .50 .56 .14 .47 .35
+CLARE N=4 .87 .84 .76 .50 .65 .50 .57 .14 .49 .37
+CLARE N=6 .87 .84 .75 .50 .63 .52 .58 .14 .48 .37
+Pixtral CLARE N=2 .89 .86 .80 .63 .58 .48 .54 .13 .49 .35
+CLARE N=4 .90 .87 .80 .62 .60 .51 .56 .14 .51 .37
+CLARE N=6 .90 .87 .80 .63 .61 .51 .55 .13 .49 .36
+Table 13: Ablation on the number of retrieved candidates (N) for classification tasks.
+Backbone Model VQA-RAD SLAKE PathVQA
+Closed Open Closed Open Closed Open
+Qwen2-vl CLARE N=2 .78 .46 .88 .82 .92 .37
+CLARE N=4 .79 .48 .90 .84 .93 .38
+CLARE N=6 .78 .47 .89 .83 .93 .38
+Pixtral CLARE N=2 .78 .45 .89 .83 .93 .38
+CLARE N=4 .78 .47 .90 .84 .93 .37
+CLARE N=6 .78 .47 .89 .84 .93 .38
+Table 14: Ablation on the number of retrieved candidates (N) for visual question answering tasks.
+D.4 CLARE Variants with Medical Pre-Trained Backbones
 Although our work focuses on general-purpose LVLMs, we also evaluated a medical-specific baseline.
@@
 Full results are presented in Table 10.
-D.4 Retrieval Robustness Analyses.
+D.5 Retrieval Robustness Analyses.
 We strive for our model to be robust in cases where the retrieval is noisy.
@@
 This could also be helpful in cases where model users would like it to be versatile and support both retrieval-augmented prediction and non-retrieval-augmented prediction (e.g., in cases where no retrieved contexts were discovered).
-As shown in Table 8 in Appendix D, removing retrieval from Jomed resulted in performance comparable or slightly lower than the reader-only model.
+As shown in Table 8 in Appendix D, removing retrieval from CLARE resulted in performance comparable or slightly lower than the reader-only model.
 Finally, we tested the model without the input image, relying only on the (image, report / caption) retrievals, to further assess the information held in retrieved data for predictions.
@@
 These results indicate that the model learns from the retrieved data, yet it does not become overly dependent on it as we previously showed.
+D.6 Different number of retrieved candidates
+Our model demonstrates robustness across varying numbers of retrieved candidates.
+We explore performance for N = 2, 4, and 6, showing consistent results across the different configurations.
+Results are presented in Table 13 for classification tasks and in Table 14 for visual question answering tasks.
 Appendix E Prompt Design
@@
 To convert open questions into closed ones with the o3 model, we use the prompt in Prompt 1.
-To assess the utility of o3 as a re-ranker of the most predictive image–caption pair (Section D.2, Appendix D), we provide both LVLM backbones (Pixtral and Qwen2-VL) with the task-specific prompt in Prompt 2.
-We also include the training prompts for Jomed across datasets—Prompts 3, 5, 4, 7, 6—and the VQA prompt (Prompt 8).
+To assess the utility of o3 as a re-ranker of the most predictive image–caption pair (Section D.3, Appendix D), we provide both LVLM backbones (Pixtral and Qwen2-VL) with the task-specific prompt in Prompt 3.
+We also include the training prompts for CLARE across datasets—Prompts 4, 6, 5, 8, 7—and the VQA prompt (Prompt 9).
 Convert the following open-ended question into a closed yes/no question based on the given answer.
@@
 Listing 1: Prompt used for o3 to convert open question into closed one
+You are an expert medical AI assistant specializing in evaluating medical image and text relevance for clinical decision-making.
+Compare two medical (image, text) pairs and determine which is MORE RELEVANT to the ground truth label.
+Ground Truth Label: "{gt_text}"
+First Image: {image_1}
+First Text: {text_1}
+Second Image: {image_2}
+Second Text: {text_2}
+Consider:
+- Medical image findings
+- Medical terminology accuracy in text
+- Clinical relevance to the ground truth
+- Diagnostic information alignment
+- How useful each (image, text) pair would be for the given condition
+Return JSON format:
+"choice": <1 or 2>,
+"confidence": <"high", "medium", or "low">,
+"explanation": "<brief 1-sentence explanation>"
+Be concise and accurate. Choose the option that is MORE medically relevant to the ground truth.
+Listing 2: Prompt used for retrieval candidate relevance evaluation
 Task:
@@
 Caption Number: [your selected number]
-Listing 2: Prompt used for o3
+Listing 3: Prompt used for o3
 […]
-Listing 3: Prompt used by Jomed for Breast classification
+Listing 4: Prompt used by Jomed for Breast classification
 […]
-Listing 4: Prompt used by Jomed for Retina classification
+Listing 5: Prompt used by Jomed for Retina classification
 […]
@@
 […]
-Listing 5: Prompt used by Jomed for Derma classification
+Listing 6: Prompt used by Jomed for Derma classification
 […]
@@
 […]
-Listing 6: Prompt used by Jomed for Vindr-PCXR classification
+Listing 7: Prompt used by Jomed for Vindr-PCXR classification
 […]
@@
 […]
-Listing 7: Prompt used by Jomed for BRSET classification
+Listing 8: Prompt used by Jomed for BRSET classification
 […]
-Listing 8: Prompt used by Jomed for visual question answering tasks
+Listing 9: Prompt used by Jomed for visual question answering tasks